### v3

# Multimodal Evaluation of Russian-language Architectures

Abstract
footnotetext: * Core contributors
Multimodal large language models (MLLMs) are currently at the center of research attention, showing rapid progress in scale and capabilities, yet their intelligence, limitations, and risks remain insufficiently understood. To address these issues, particularly in the context of the Russian language, where no multimodal benchmarks currently exist, we introduce MERA Multi, an open multimodal evaluation framework for Russian-spoken architectures. The benchmark is instruction-based and encompasses default text, image, audio, and video modalities, comprising 18 newly constructed evaluation tasks for both general-purpose models and modality-specific architectures (image-to-text, video-to-text, and audio-to-text). Our contributions include: (i) a universal taxonomy of multimodal abilities; (ii) 18 datasets created entirely from scratch with attention to Russian cultural and linguistic specificity, unified prompts, and metrics; (iii) baseline results for both closed-source and open-source models; (iv) a methodology for preventing benchmark leakage, including watermarking for private sets. While our current focus is on Russian, the proposed benchmark provides a replicable methodology for constructing multimodal benchmarks in typologically diverse languages, particularly within the Slavic language family.
Multimodal Evaluation of Russian-language Architectures
Artem Chervyakov*1, Ulyana Isaeva*1, Anton Emelyanov1, Artem Safin1, Maria Tikhonova1,4, Alexander Kharitonov1, Yulia Lyakh1, Petr Surovtsev1, Denis Shevelev1, Vildan Saburov1,8, Vasily Konovalov3,5,6, Elisei Rykov5,7, Ivan Sviridov2, Amina Miftakhova2, Ilseyar Alimova5, Alexander Panchenko5,3, Alexander Kapitanov1, Alena Fenogenova*1 1SberAI, 2Sber AI Lab, 3AIRI, 4HSE University, 5Skoltech, 6MIRAI, 7T-Tech, 8Moscow Center for Advanced Studies Correspondence: mera@a-ai.ru
Figure 1: Overview of MERA Multi. The benchmark unites multimodal evaluation, taxonomy-based skill assessment, and data leakage protection across 18 tasks covering (default) text, image, audio, and video modalities. It employs standardized block-prompting, compound scoring, and integrates methods for multimodal content protection, forming a transparent and robust methodology for culturally grounded multimodal evaluation in Russian.
1 Introduction
Recent breakthroughs in generative AI, including models like GPT-5111https://openai.com/index/introducing-gpt-5, ImageBind Girdhar et al. (2023), and LLaVa Liu et al. (2024a), have significantly advanced the state of the art across multiple modalities. This accelerated progress created a growing need for comprehensive multimodal benchmarks capable of rigorous evaluation of the full spectrum of versatile capabilities of such models. Although several benchmarks have been proposed for English and general-domain evaluation, such as MultiBench Liang et al. (2021), MMBench Liu et al. (2024e), and General-Bench Fei et al. (2025), they predominantly neglect the linguistic and cultural nuances of Slavic languages, particularly Russian. Beyond its Cyrillic script, Russian possesses a rich cultural context where concepts familiar to native speakers (e.g., folklore, Soviet media) are foreign to others, creating a challenge for automatic understanding.
Existing Russian-specific benchmarks, including TAPE Taktasheva et al. (2022), Russian SuperGLUE Fenogenova et al. (2022), and MERA Fenogenova et al. (2024), focus exclusively on text-based tasks, leaving a critical gap in multimodal evaluation. To address this, we introduce MERA Multi [^2] https://mera.a-ai.ru/en/multi, the first multimodal benchmark for MLLM evaluation in Russian. It comprises 18 tasks spanning (default) text [^3] Since textual input is an inherent component of all tasks in our benchmark, we treat it as the default modality. Consequently, we do not explicitly list it when describing or tabulating tasks, unless required for clarity., image, audio, and video modalities, built upon a unified taxonomy of multimodal abilities. Beyond Russian MERA Multi offers a blueprint for developing multimodal benchmarks across other Slavic and morphologically rich languages. These languages share structural complexity, typological proximity, and cultural specificity that make direct translation or adaptation of English-centric benchmarks inadequate. Thus, it serves not only as the first multimodal benchmark for Russian, but also as a scalable methodology for culturally and linguistically aware evaluation in underrepresented settings.
More specifically, our contributions are fourfold:
  • We propose a unified taxonomy and evaluation methodology designed for MLLMs assessment;
  • We create 18 novel datasets  [^4] https://hf.co/collections/MERA-evaluation/mera-multimodality incorporating Russian cultural and linguistic specificities, unified prompts, and metrics.
  • We provide baseline performance results for both open- and closed-source models;
  • We establish a data leakage analysis and watermarking strategy to protect private evaluation datasets.
Additionally, we provide a standardized codebase for full reproducibility and a submission platform with automated scoring and a public leaderboard [^5] https://mera.a-ai.ru/en/multi/leaderboard. All datasets and code are made available under the MERA Multi license, which allows the use of the benchmark sets for non-commercial purposes, provided that the data is not used for model training of any kind. These elements establish a foundation for transparent and culturally aware multimodal evaluation in Russian, while positioning MERA Multi as a reference point for developing similar frameworks across Slavic and other non-English languages, thereby promoting broader community development and cross-linguistic benchmarking.
Benchmark Modalities # Tasks/Skills Primary Focus Language Cultural Focus
MultiBench Liang et al. (2021) 10 (Text, Image, Audio, Video, Table, Set, …) 20 General-purpose representation EN General
MMBench Liu et al. (2024e) Image, Text 20 Fine-grained visual reasoning EN, ZH General
SUPERB Yang et al. (2021) Speech, Audio 10 Speech processing EN General
STAR Wu et al. (2021) Video, Text 4 VideoQA, temporal reasoning EN General
InfiniBench Ataallah et al. (2024) Video, Text 8 Long-context video understanding EN General
Video-MME Fu et al. (2025) Video, Text 12 Fine-grained video analysis EN General
General-Bench Fei et al. (2025) Text, Image, Video, Audio, Tabule 700+ Large-scale ability coverage EN General
OmniDialog Razzhigaev et al. (2024) Text, Image, Audio 8 Multimodal dialogue EN General
MMMU Yue et al. (2024) Image, Text 30 Expert-level reasoning EN General
SEED-Bench Li et al. (2023) Image, Video, Text 12 Generative comprehension EN General
MERA Multi (ours) Text, Image, Audio, Video 18 Comprehensive understanding RU General + Russian
Table 1: Comparison of major multimodal evaluation benchmarks and MERA Multi.
2 Related Work
Text-Based Benchmarks
Evaluation of language models has historically relied on Natural Language Understanding benchmarks such as GLUE Wang et al. (2019b) and SuperGLUE Wang et al. (2019a), which set initial standards for English. As these benchmarks became saturated, newer instruction-oriented and reasoning-focused benchmarks emerged such as BIG-bench Suzgun et al. (2023) and HELM Perlitz et al. (2024). Further efforts such as MMLU Hendrycks et al. (2021), AGIEval Zhong et al. (2024), and C-Eval Huang et al. (2023) extended evaluation to academic and professional domains, though primarily in English and Chinese. For Russian, several text-based benchmarks were introduced. Among them are Russian SuperGLUE Shavrina et al. (2020); Fenogenova et al. (2022), TAPE Taktasheva et al. (2022), and RuCoLA Mikhailov et al. (2022). The instruction-based benchmark MERA Fenogenova et al. (2024) advanced this line of work but remained limited to text-only evaluation. Our work extends this direction toward multimodal, instruction-following assessment of MLLMs in Russian.
Multimodal Benchmarks
The rapid progress of multimodal models has led to numerous benchmarks extending evaluation beyond text. Early large-scale efforts as MultiBench Liang et al. (2021) covered 10 modalities and emphasized general-purpose representation learning. MMBench Liu et al. (2024e) focused on fine-grained visual reasoning and bilingual (English–Chinese) evaluation, while SUPERB Yang et al. (2021) unified diverse audio tasks under a single framework. For video understanding, STAR Wu et al. (2021), InfiniBench Ataallah et al. (2024), and Video-MME Fu et al. (2025) advanced evaluation toward temporal and long-context reasoning. More comprehensive setups such as General-Bench Fei et al. (2025), OmniDialog Razzhigaev et al. (2024), MMMU Yue et al. (2024), and SEED-Bench Li et al. (2023) assess multimodal and reasoning abilities at scale. All in all, despite these advances, most existing multimodal benchmarks are English-centric. To our knowledge, there is currently no existing multimodal benchmark for the Russian language. MERA Multi addresses the gap by providing a unified and culturally adapted evaluation of understanding across several modalities (see Table 1 for the detailed comparison between MERA Multi and other benchmarks).
3 Overview of MERA Multi Benchmark
Dataset / task Size HB Answer License
audio ruEnvAQA 596 0.95 MC CC BY-NC 4.0
RuSLUn 741 0.90 OE $C$-BY-4.0
*AQUARIA 738 0.98 MC MERA Multi
*ruTiE-Audio 1500 0.75 MC MERA Multi
image ruCLEVR 1148 0.96 OE $C$-BY-4.0
ruCommonVQA 3015 0.84 OE $C$-BY-4.0
ruNaturalScienceVQA 363 0.99 MC $C$-BY-SA-4.0
WEIRD 814 0.85 MC $C$-BY-4.0
*LabTabVQA 339 0.91 MC MERA Multi
*RealVQA 773 0.63 OE MERA Multi
*ruHHH-Image 595 0.89 MC MERA Multi
*ruMathVQA 502 0.95 OE MERA Multi
*ruTiE-Image 1500 0.77 MC MERA Multi
*SchoolScienceVQA 4227 0.82 MC MERA Multi
*UniScienceVQA 7432 0.13 OE MERA Multi
video CommonVideoQA 1200 0.96 MC $C$-BY-4.0
*RealVideoQA 671 0.96 MC MERA Multi
*ruHHH-Video 911 0.84 MC MERA Multi
Table 2: Overview of datasets in MERA Multi. Those marked with an asterisk are private datasets collected from scratch, while the others are public datasets compiled from open-source datasets. Size column shows the number of samples in the dataset. HB is the human baseline value (basic / expert or basic only, see F.2 for details). Answer column is the task format (MC and OE stand for multiple-choice and open-ended, respectively). MERA Multi license refers to the benchmark license anonymized for the review period.
Figure 1 presents an overview of MERA Multi and its methodological parts.
3.1 Benchmark General Structure
The proposed benchmark is designed to evaluate the capabilities of Russian MLLMs. Besides the omnipresent textual modality, it incorporates tasks from three other modalities: image (11 datasets), audio (4 datasets), and video (3 datasets). The tasks are formulated in two primary formats: multiple-choice questions and open-ended questions requiring short free-form answers.
To balance reproducibility and novelty, the benchmark integrates both publicly available datasets (7 tasks), curated from open data sources, and private datasets (11 tasks), collected from scratch specifically for this study. The latter are designed to incorporate Russian cultural nuances, target underexplored skill categories, and mitigate potential data contamination issues (see details in section 3.4). A complete list of datasets is provided in Table 2. Task examples and dataset creation details are in A.
3.2 Skill Taxonomy
Contemporary studies on multimodal benchmarks often rely on custom skill sets during the dataset design process Liu et al. (2024e); Fu et al. (2024). We performed a comprehensive analysis of such systems and synthesized them into a consolidated MLLM skill taxonomy, which underpins the foundation of MERA Multi. Such a system functions as a comprehensive map for skill coverage, which, when coupled with an alignment of existing datasets to corresponding skills, can spotlight deficiencies in benchmark task diversity.
The taxonomy is aligned with three broad categories of human-like cognition: knowledge, perception, and reasoning, a division also adopted by recent multimodal benchmarks. The knowledge taxonomy is shown in Table 3, and the perception and reasoning parts are provided in the Appendix in Table 7 and Table 8 respectively, due to space constraints. To comply with the emerging MLLM capabilities, this taxonomy is designed to be extendable and this paper provides the initial version of the unified taxonomy and encourages its adoption and extension in future research.
Each MERA Multi dataset is systematically mapped to a predefined set of multimodal skills. Rather than assigning a single skill per task, each task is designed to be multifaceted, evaluating a distinctive combination of abilities. For example, a single task might require visual perception (to identify objects), OCR (to read text in the image), and reasoning (to answer a why-question) all together.
Further details on skill taxonomy are in Appendix B.
Taxonomy Level Modality
L1 L2 L3 L4 L5 Image Audio Video
Knowledge Knowledge Common everyday knowledge Common everyday knowledge ruCommonVQA
RealVQA ruTiE-Audio CommonVideoQA
ruHHH-Image ruEnvAQA RealVideoQA
WEIRD AQUARIA ruHHH-Video
ruTiE-Image
Ethics ruHHH-Image ruHHH-Video
Domain knowledge Common domain knowledge ruMathVQA ruTiE-Audio CommonVideoQA
RealVQA ruEnvAQA RealVideoQA
ruTiE-Image AQUARIA
Expert domain knowledge ruMathVQA
ruNaturalScienceVQA
UniScienceVQA
SchoolScienceVQA
Table 3: Knowledge taxonomy structure and multimodal task distribution in MERA Multi. Columns L1-L5 show the skill hierarchy levels, while Image, Audio and Video columns indicate which tasks cover each knowledge type across different modalities.
3.3 Evaluation Methodology
MERA Multi evaluation methodology is designed to systematically assess the multimodal reasoning, perception, and knowledge abilities of MLLMs. It measures both general-purpose and modality-specific competence across image, audio, and video inputs. To achieve this, we propose the benchmark that integrates four complementary components: (i) a block-prompting structure that ensures consistency and diversity of task formulations across modalities (Section 3.3.2); (ii) dual-level evaluation metrics combining symbolic and semantic correctness through a dedicated LLM-as-a-judge model (Section 3.3.3); (iii) an evaluation pipeline including scoring aggregation and cross-modality weighting (Section 3.3.4); and (iv) a submission protocol enabling automated, reproducible leaderboard updates (Section 3.3.5). These components provide a coherent methodology that balances rigor, interpretability, and cross-model comparability in multimodal evaluation.
3.3.1 Evaluation Framework Implementation
We build the benchmark code base on the lm-eval [^6] https://github.com/EleutherAI/lm-evaluation-harness framework (Gao et al., 2024; Biderman et al., 2024), extending it to support multimodal inputs while preserving its core structure for texts. The codebase introduces comprehensive vision, audio, and video evaluation through, and chat-template formatting tailored for instruction-tuned and API-based models. Multimodal data is integrated either by passing it separately to the model’s processor or by embedding it directly within a chat template alongside the text. All evaluations are strictly generative: models produce free-form answers until a stop condition is met. All outputs are used directly except RuSLUn which requires minimal parsing for structured output.
To balance rigor and semantic understanding, we employ two primary metrics across most datasets. ExactMatchExactMatch (EMEM) serves as a generative analog of accuracy, using normalized string comparison to assess both factual correctness and strict adherence to the specified output format. Complementing this, the JudgeScoreJudgeScore (JSJS) measures semantic similarity via an LLM-as-a-judge (trained for this purpose) that scores an answer as 1 for substantive agreement with the reference and 0 otherwise. This dual approach provides a nuanced view: EM is sensitive to format, while the JS focuses on meaning. For the datasets where this format is unsuitable we employ task-specific metrics or report only the JS.
The $FinalScore$ of the task (for Total Score calculation purposes) is defined as follows:
$FS(task)=\frac{\text{EM}(task)+\text{JS}(task)}{2};$ (1)
It should be noted that as long as some of our prompts suggest that the model return the answer for the question after a specific phrase (“ANSWER”), we consider $EM$ to be the maximum of two scores: (i) $EM$ of the full model answer, (ii) $EM$ of the last part of the model’s answer split by the previously mentioned specific phrase. If the model breaches the instruction, we assess the full answer. Otherwise, we consider the part that is after the specific phrase to be the model’s answer as required by the instruction.
3.3.2 Prompt Structure
Following the approach of Voronov et al. (2024), we avoid hard-coding task-specific prompts by employing a block-prompting scheme that constructs universal prompts aligned with each task’s structure. For every dataset, we instantiate a predefined set of 10 prompts drawn from fixed block layouts that we initially designed and realized under different surface modalities (e.g., formal request, command, technical statement). This design preserves a uniform formulation across benchmark tasks while allowing for controlled variation, such as the presence or absence of a reasoning request, to mitigate benchmark saturation.
Following the methodology of Fenogenova et al. (2024), we uniformly assign prompts for each task across dataset samples to ensure that the aggregated performance metric reflects an average over prompt variants rather than the idiosyncrasies of a single prompt. A post-hoc statistical analysis (see Appendix E for the details) confirmed that our prompt set is not invariant, as some prompts produce statistically significant shifts in model scores. This finding directly motivates and validates our multi-prompt design as a necessary guard against evaluation bias.
3.3.3 Judge Scoring
Evaluating open-ended model generations requires moving beyond strict heuristics. We frame this as a task of assessing semantic equivalence, conditioned on the question, as devising universal regex patterns to extract answers from free-form text, particularly from multi-step reasoning chains – proves infeasible. The practically infinite space of possible valid generations necessitates a more flexible and scalable assessment approach. We therefore introduce a learned LLM-as-judge model, framing evaluation as a question-conditioned semantic equivalence task (binary classification – determine the correctness of a model’s prediction).
This judge is trained on a diverse, human-annotated dataset of model outputs collected from various Russian-language instruction-based benchmarks. The final model, based on the RuModernBERT [^7] https://hf.co/deepvk/RuModernBERT-base Spirin et al. (2025) encoder, achieves F1 score of 0.96 on a held-out test set and shows 99.6% agreement with EM on identical answers, confirming its reliability. We deploy this judge as our primary metric, providing a robust, scalable score for correctness across the benchmark. Comprehensive details on data, training, and model comparisons are in the Appendix D.
3.3.4 Evaluation Pipeline
Let MERA Multi comprise $M$ modalities (e.g., image, audio, video) not counting the omnipresent text one, with modality $m$ containing $n_{m}$ tasks. We treat modalities as equally important and split each modality’s weight evenly across its tasks. Let $k_{m}$ be the number of tasks a model attempted in modality $m$ with $s_{m,i}\in[0;1]$ being the score (average of all task $i$ metrics).
  • The Attempted Score. Quality over the tasks the model attempted, with the original equal-per-modality weighting renormalized over the attempted subset:
    $A=\frac{\sum_{m=1}^{M}\frac{1}{n_{m}}\sum_{i\in A_{m}}s_{m,i}}{\sum_{m=1}^{M}\frac{k_{m}}{n_{m}}}$ (2)
  • The Coverage. Breadth of evaluation is the fraction of the benchmark the model actually attempted:
    $C=\frac{1}{M}\sum_{m=1}^{M}\frac{k_{m}}{n_{m}}$ (3)
  • The Total Score. We combine quality and breadth as
    $T=A\cdot C$ (4)
This approach has the following positive effects:
  • Separation of concerns. $A$ reports how well a model performs where it runs; $C$ reports how much of MERA Multi it covers.
  • Fair, single leaderboard. $T$ enables joint ranking without imputing arbitrary zeros for missing modalities; specialists still excel on modality-specific boards (higher $A$, lower $C$).
  • Stable under growth. Adding/removing tasks adjusts $C$ (breadth), but leaves $A$ (quality on attempted tasks) invariant; equal-per-modality weighting prevents any modality from dominating due to task count.
3.3.5 Submission
The private dataset test answers are available only for the organizers and experts supporting the benchmark. The scoring system is automatic and is available on the benchmark platform.
The process of submission is the following:
  • First, users clone MERA Multi benchmark repository and form submission files using shell script and the provided code.
  • Second, they upload the submission files via the platform interface [^8] For this step the registration on the platform is required. in their personal account for automatic assessment.
  • The evaluation result is then displayed in the user’s account and kept private unless they request publication using the “Publish” function.
  • In this case, it undergoes an expert verification of its reproducibility, which includes checking log files automatically formed by the evaluation script and the provided submission information. Once approved, the model’s score is shown publicly on the leaderboard (user can choose on which leaderbord image/audio/video/multi to add the results), while its specific outputs remain private.
3.4 Data Protection
As pre-training datasets grow, benchmark leakage is becoming more common. This issue is exacerbated by opaque training processes and the undisclosed use of supervised data in modern MLLMs, which undermines the validity of benchmarks and fosters unfair comparisons, ultimately slowing progress. To protect the multimodal data in our private benchmark suite from unauthorized use, we employ: (i) data watermarking (Section 3.4.1), (ii) leakage detection (Section 3.4.2). We also introduce the license, which explicitly permits the use of the benchmark data for research and non-commercial testing purposes, but strictly prohibits its incorporation into any model training process. Together, these measures provide robust technical and legal safeguards.
3.4.1 Data Watermarking
We embed imperceptible yet identifiable watermarks into our benchmark data to trace its provenance and detect unauthorized use in training corpora. Our approach is tailored to each modality:
  • Audio: We use AudioSeal Roman et al. (2024) for localized detection, which employs a neural, inaudible watermark.
  • Images/Video: a simple overlay of the MERA Multi watermark on images and every video frame (same code across frames).
Model Confidence Interval
image Qwen2-VL-2B-Instruct (0, 0.352)
Qwen2-VL-7B-Instruct (0, 3.920)
Qwen2.5-VL-3B-Instruct (0.259, 2.180)
Qwen2.5-VL-7B-Instruct (0.444, 1.389)
llava-1.5-7b-hf (0, 0.937)
llava-1.5-13b-hf (0, 0.839)
llama3-llava-next-8b-hf (0, 1.024)
gemma-3-12b-it (1.183, 3.120)
gemma-3-12b-it (0, 1.756)
video LLaVA-NeXT-Video (0.811, 3.314)
LLaVA-NeXT-Video-DPO (1.367, 2.125)
LLaVA-NeXT-Video-34B (0, 1.373)
Qwen2-VL-2B-Instruct (0, 0.645)
Qwen2.5-VL-7B-Instruct (0, 0.781)
Qwen2.5-VL-3B-Instruct (0, 0.330)
Qwen2-VL-7B-Instruct (0, 0.105)
audio ultravox-v0_2 (0, 0.940)
ultravox-v0_3-llama-3_2-1b (0, 1.566)
ultravox-v0_5-llama-3_2-1b (0, 0.195)
ultravox-v0_6-llama-3_1-8b (0, 0.875)
ultravox-v0_4 (0, 0.401)
ultravox-v0_4_1-mistral-nemo (0, 0.079)
Qwen2-Audio-7B-Instruct (0.0, 0.0)
MiniCPM-o-2_6 (0.0, 0.0)
Table 4: Confidence intervals of Judge Score (JS) differences between watermarked and clear data (%).
Table 4 demonstrates the Confidential Intervals (CI) for differences of the models’ metrics on data with watermarks and without them. The CI are computed for differences in per cents for convenience. The results demonstrate that with 95% probability the difference is less than 5% (usually less than 1%) which leads to the conclusion: our watermarking strategy does not significantly affect the evaluation results.
3.4.2 Data Leakage Detection
We detect training-set data leakage using membership inference attacks Shokri et al. (2017). Our approach Emelyanov et al. (2025) extends the Semantic Membership Inference Attack (SMIA) Mozaffari and Marathe (2024) to MLLMs, calculating the loss for a data point by considering its text in conjunction with its paired image, video, or audio data. The Multimodal SMIA (MSMIA) method identifies leakage by comparing a model’s behavior on original examples versus their semantically perturbed "neighbors" (masked, removed, doubled, switched text tokens); models that have been trained on the data (models with data leak) are to exhibit systematically different loss patterns.
Modality AUC-ROC
Image 88.658
Video 88.388
Audio 81.250
Table 5: Average AUC-ROC of MSMIA per modality. Averaging over the models used for training and evaluating MSMIA.
Concretely, we train the MSMIA detector by comparing two versions of a model: the original and a version we fine-tune on candidate data (simulating a leak). The detector learns to distinguish between them by analyzing the differences in loss and text embeddings when processing original data points versus their perturbed neighbors. Once trained, this detector can be applied to any other model to output a probability that a specific data sample was part of that model’s training set. Overall results of the MSMIA detection capabilities are presented in Table 5. Following the original methodology, we evaluate detection performance using AUC-ROC. The table demonstrates relatively high scores, which means that the MSMIA method tends to be capable of detecting whether a model has been trained on some multimodal data samples or not, with high probability and a rather low false-positive rate. The details on the MSMIA training and the metrics are provided in Appendix C.
4 Baselines
4.1 Model Baselines
We evaluate over 50 publicly available multimodal models from the most trending model families on HuggingFace, varying in size from 1B to 110B parameters. Also we evaluate proprietary GPT 4.1 (OpenAI) to make a comparison between open and closed source models.
See Appendix F.1 for the baseline details. We evaluate models in the same environments and scenarios by the procedure described in Section 3.3.4 and the submission procedure described in Section 3.3.5. We also provide examples of particular submissions (the model evaluated on part of the tasks of a modality).
4.2 Human Baselines
To estimate human-level performance across MERA Multi tasks, we compute human baseline (HB) values based on the aggregated results of crowd annotators. For datasets requiring additional domain expertise, we also establish expert HB obtained from qualified expert annotators. Annotation quality is ensured through honeypot tasks with automated correctness verification and post-hoc filtering of low-performing annotators. All crowd annotations are collected via the ABC Elementary platform [^9] https://app.elementary.center, which guarantees data anonymity, fair compensation, and ethical compliance. See Appendix F.2 for detailed methodology and cost breakdown .
Model Total Score Attempted Score Coverage Image Total Audio Total Video Total
Qwen3-Omni-30B-A3B-Inst 0.500 0.563 0.889 0.554 0.561 0.410
Qwen2.5-Omni-7B 0.317 0.317 1.000 0.226 0.474 0.442
Qwen2.5-VL-72B-Inst 0.302 0.453 0.667 0.406 0.000 0.625
MiniCPM-o-2_6 0.255 0.255 1.000 0.182 0.369 0.373
Qwen2-VL-72B-Inst 0.254 0.381 0.667 0.333 0.000 0.557
Qwen2.5-Omni-3B 0.251 0.251 1.000 0.180 0.380 0.337
Qwen2.5-VL-7B-Inst 0.209 0.313 0.667 0.256 0.000 0.523
GPT 4.1 0.159 0.478 0.333 0.478 0.000 0.000
Qwen2-VL-7B-Inst 0.145 0.218 0.667 0.195 0.000 0.301
Qwen2.5-VL-3B-Inst 0.136 0.203 0.667 0.142 0.000 0.427
InternVL3-9B 0.135 0.203 0.667 0.172 0.000 0.316
Qwen3-VL-2B-Inst 0.125 0.187 0.667 0.125 0.000 0.416
ultravox-v0_5-llama-3_1-8b 0.104 0.311 0.333 0.000 0.311 0.000
ultravox-v0_4_1-llama-3_1-8b 0.102 0.307 0.333 0.000 0.307 0.000
ultravox-v0_4 0.101 0.304 0.333 0.000 0.304 0.000
ultravox-v0_6-llama-3_1-8b 0.100 0.300 0.333 0.000 0.300 0.000
Phi-4-multimodal-instruct 0.098 0.147 0.667 0.185 0.042 0.000
ultravox-v0_4_1-mistral-nemo 0.088 0.265 0.333 0.000 0.265 0.000
audio-flamingo-3-hf 0.086 0.259 0.333 0.000 0.259 0.000
llava-next-110b-hf 0.079 0.236 0.333 0.236 0.000 0.000
Phi-3.5-vision-inst 0.076 0.228 0.333 0.228 0.000 0.000
Qwen2-Audio-7B-Inst 0.074 0.223 0.333 0.000 0.223 0.000
SmolVLM-Inst 0.064 0.192 0.333 0.192 0.000 0.000
gemma-3-27b-it 0.050 0.151 0.333 0.151 0.000 0.000
granite-vision-3.3-2b 0.048 0.143 0.333 0.143 0.000 0.000
deepseek-vl2-small 0.044 0.163 0.273 0.133 0.000 0.000
Table 6: Overall baselines information over three modalities (vision/image, audio, video). All scores are aggregated. Modality total score is the attempted score multiplied by coverage of the modality.
5 Results
The leaderboard is designed in such a way that the more modalities the model covers, the higher the Total Score could be. The top performer, Qwen3-Omni-30B-A3B-Instruct, leads with a Total Score of 0.5, driven by its high Attempted Score (0.563) and rather high Coverage (0.889), showing strong image, audio, and video capabilities. Notably, the models from Qwen families obtain larger scores for image and video modalities (first 3 places of the overall leaderboard are taken by those models). GPT 4.1 still leads in image modality while having low Coverage (0.333), which leads to a lower Total Score (0.159 compared to 0.5 of the top performer). The main trend is defined by the metrics used: broader coverage leads to higher Total Score. Thus omni-models occupy the first places. But strong uni- or bimodal capabilities may gain advantage over middle-performing models with high Coverage (e.g. Qwen2.5-VL-72B-Instruct and Qwen2-VL-72B-Instruct with 0.302 and 0.254 Total Scores respectively).
This pattern is consistent across all modalities. In audio, the specialists from ultravox family tend to display poorer performance compared to omni-models like Qwen2.5-Omni-7B (0.311 vs 0.474 for Audio Total Score) even though ultravox models use other LLM’s from Mistral, Llama, Qwen families as backbones. Considering the video modality, vision models are usually trained with video-inputs or may slice the video into frames and use “regular” vision encoders for them, which explains why Qwen2-VL-72B-Instruct shows the best Video Total Score (0.625) while the models that specialize specifically on video modality like those from LLaVA-NeXT-Video family show poorer metrics.
Consistently, Judge Score (JS) >> EM across models, indicating that many responses are semantically correct but violate output format; whenever JS $\approx$ EM, the model followed instructions closely. This gap justifies reporting JS alongside EM. When JS << EM, this means that we can extract the answer from model’s generation but the entire generation may be misleading (e.g. wrong rationale, reasoning conclusion mismatches the selected answer).
Tables with separate datasets metrics and analysis may be found in Appendix F.1.
Takeaway 1: There is still a gap between modalities. Omni-models partially bridge it. Specialist models, however, show that while image understanding is a relatively mature field, audio and video understanding are underrepresented in terms of both models and datasets, which is reflected in lower scores on benchmark tasks.
Takeaway 2: Overall metrics are robust to missing task scores (unfinished submission) and multiple modalities.
Conclusion
The rapid progress of generative AI has introduced new challenges for evaluating models in multimodal contexts. We present MERA Multi, the first comprehensive framework for transparent and culturally grounded multimodal evaluation in Russian. The benchmark encompasses 18 tasks across four modalities (default text, image, audio, and video), covering diverse domains and scenarios. It systematizes diverse multimodal abilities via a proposed taxonomy and evaluates them through methodologically verified prompts and metrics. We also provide a standardized code base [^10] https://github.com/MERA-Evaluation/MERA_MULTIMODAL that guarantees reproducibility and a submission platform offering automated evaluation, scoring, and open leaderboards.
In the future, we plan to expand the benchmark to encompass additional scenarios and actively encourage community contributions. We envision MERA Multi as a collaborative initiative that promotes transparent evaluation practices and provides a methodological foundation for developing culturally aware multimodal benchmarks across non-English languages such as Slavic, ultimately advancing the creation of more robust and reliable multimodal models.
Limitations
First of all, despite the fact that our benchmark covers 18 tasks spanning multiple domains, aiming to represent complementary semantic abilities of the models, this set may be underrepresenting some abilities of the model or some domains which may be crucial for certain tasks and applications. Namely, it is not impossible that a model which excels at our benchmark will perform poorly on a specialized domain or task.
Even with fixed prompts and decoding settings, MERA Multi scores can vary because the entire hardware–software stack affects inference: GPU model, drivers/CUDA/cuDNN, PyTorch, vLLM/transformers (and commit hashes), flash-attention kernels, tokenizers/checkpoints, precision/quantization, and batching — some of which are non-deterministic. We therefore request public submissions adhere to the same parameters and, in submission information, specify the GPUs and libraries versions they used for reproducibility purposes.
Ethical Statement
While the presented benchmark is able to comprehensively evaluate “semantic” abilities of the model, i.e., the capacity of individual models to understand and reason about data in different modalities, we did not perform explicit evaluation of any bias of these models, e.g., toward any kind of underrepresented minorities. In our opinion, this is an extremely important direction of future work, yet being outside the scope of our current contribution.
For the creation of novel datasets, the work of human annotators is used. We state that their work was adequately paid or compensated (see Section F.2 for the details).
Researchers participating in the benchmark will be encouraged to adhere to ethical research practices, including proper citation, acknowledgment of data sources, and responsible reporting of results. Regular ethical reviews will assess the benchmark’s impact, identify potential ethical concerns, and implement necessary adjustments to uphold the highest ethical standards throughout the usage usage.
We proofread the text of this article using Overleaf Writefull assistant [^11] https://www.writefull.com, GPT-4o [^12] https://chatgpt.com, Grammarly [^13] https://app.grammarly.com to correct grammatical, spelling, and style errors and paraphrase sentences. We emphasize that these tools are used solely to enhance the quality of English writing, in full compliance with the ACL policies on the responsible use of AI writing assistance. Nevertheless, some segments of our publication can be potentially detected as AI-generated, AI-edited, or human-AI-generated.
Acknowledgments
MERA is a collaborative project, thoughtfully designed to serve the needs of both industry and academia. The authors extend their sincere gratitude to our partners from the AI Alliance Russia, whose invaluable collaboration made an undertaking of this scale possible. Special thanks are due to Ekaterina Morgunova, Yegor Nizamov, and Uliana Plugina for their significant contributions in coordinating our benchmark partners and contractors for the website development.
We also express our deep appreciation to the entire team dedicated to developing the website platform and maintaining the scoring system.
We are profoundly grateful to the following individuals for their dedication and hard work: Yaroslav Grebnyak, Anna Kostikova, Aleksandr Sautin, Artem Goryainov, Aleksandra Rak, Artem Goryainov, Albina Akhmetgareeva, Igor Churin, Leonid Sinev, Yulia Lazareva, Ksenia Biryukova, Jamilya Erkenova, Valentina Khlebutina, Maria Slabunova, Sergei Markov and to many others whom we may have inadvertently missed, but who supported us with their ideas, collaborated on dataset creation, validated results, and helped organize the workflow — your contributions are sincerely appreciated.
We also thank Zaryana Damashova and Ekaterina Artemova for their contributions to the creation of the RuSLUn dataset.
The work of Alexander Panchenko was supported by the RSF project 25-71-30008 “Laboratory for reliable, adaptive, and trustworthy Artificial Intelligence”.
Appendix
Appendix A Dataset Description
A.1 AQUARIA
The dataset includes multiple-choice questions that test complex audio comprehension, including speech, non-verbal sounds, and music. The tasks in the dataset require not only the recognition of speech but also the analysis of the entire auditory situation and the interactions among its components. The audio tracks used in AQUARIA were created specifically for this dataset.
The dataset contains 9 types of tasks:
  • Audio scene classification
  • Audio captioning (matching audio with its textual description)
  • Audio comparison (finding differences between two audio)
  • Audio sequence analysis
  • Emotion recognition (recognition of emotions and subjective characteristics of a speaker)
  • Sound QA (questions related to analysis of non-verbal signals)
  • Speaker characterization (recognition of objective characteristics of a speaker)
  • Music QA (questions requiring analysis of music and related knowledge)
  • Music characterization (recognition of objective characteristics of music)
Question. What is the difference between the two provided audio recordings? Audio_1. samples/audio194.wav Audio_2. samples/audio195.wav A. In the first recording, a door is being unlocked; in the second, it was already unlocked B. In the first recording, the door creaks; in the second, it doesnt C. In the first recording, a woman enters the apartment; in the second - a man D. In the first recording, a person enters through an open door; in the second, they unlock the lock Answer. B
Motivation
The methodology for evaluating large audio-language models (LALMs), as well as the models themselves, is a fairly recent area of research. Compared to benchmarks in the vision-language domain, there are significantly fewer comprehensive benchmarks available for evaluating audio-language models. Examples of such benchmarks include AIR-Bench Yang et al. (2024), AudioBench Wang et al. (2025a), and MMAU Sakshi et al. (2025). Audio understanding tasks are generally classified into three categories: speech analysis, non-verbal signal analysis, and music analysis.
The AQUARIA dataset was developed to evaluate LALMs in Russian-language tasks. The model needs to be able to process audio because answering questions requires analyzing the associated audio track. The dataset contains 9 question types, which vary both by task category and by the model abilities they test. The dataset assesses three skill categories for audio-language models: perception, knowledge, and reasoning.
Dataset creation
Based on an analysis of existing benchmarks for testing language models with audio interfaces, we have developed 9 types of tasks that evaluate various groups of skills for these models. For each task type, experts created scenarios with dialogues, background sounds, and music, along with corresponding questions tailored to different task formulations. All scenarios were recorded using professional studio recording equipment, with voluntary use of dataset contributors’ voices. For some of the Music QA and Music characterization questions, the music tracks were created using generative models (including suno.com).
A.2 CommonVideoQA
CommonVideoQA is a public Russian-language question-answering dataset designed for evaluating video-text models (Video-LLMs), comprising questions related to video clips. It comprehensively assesses the following competencies: general video comprehension and detail recognition, possession of common and domain-specific knowledge, ability to determine the precise order of actions within a video and reconstruct the complete sequence, capability to count objects and actions over time, as well as the skill to associate actions with corresponding temporal boundaries in the video. Given an input video and a question, the task requires selecting the single correct answer from four provided options. Correct answers do not require audio track comprehension. All video clips are sourced from open public repositories.
Question. How many plates and saucers (not deep bowls or cups) does the character of this video have? A. Fifteen. B. Thirteen. C. Twelve. D. Sixteen. Answer. A
Motivation
Most published benchmarks in video understanding focus on English-language content, and currently no Russian-language benchmark is available in the public domain. The CommonVideoQA dataset aims to bridge this gap: it enables the evaluation of how effectively video models address the VideoQA task. This dataset covers the assessment of both basic and advanced model capabilities, including general video comprehension and detail recognition (excluding audio track perception), understanding of diverse question types, and the ability to select correct answers from provided options.
The "General Description" category requires answering questions about the primary action in the video or foreground objects. Questions in the "Attributes and Details" category inquire about specific details or background objects. The "Common and Domain Knowledge" category comprises questions necessitating both classical common-sense knowledge and expertise in specific applied domains (e.g., "In what order should the presented dish be prepared?"). The "Action Sequences" category includes questions testing the understanding of actions in the video, their sequential order, and the ability to reconstruct this sequence. The "Counting" category involves questions assessing the capability to count objects, repetitions of actions over time, and perform basic arithmetic operations with the counts. The "Temporal Intervals" category evaluates the ability to associate actions with temporal boundaries (video timestamps) during which these actions occur. Thus, the dataset evaluates key competencies essential for the video domain.
The dataset comprises video scenes spanning the following domains: "kitchens" (encompassing household activities), "sport" (involving training sessions or competitions), "flora and fauna" (featuring landscapes, wildlife, or plants), "tools" (demonstrating the use of various implements or auxiliary items), and "hobbies" (covering a range of personal pursuits). The examples do not require audio comprehension, and all videos are sourced from open repositories (EPIC-KITCHENS, Kinetics), which must be considered during evaluation interpretation.
Dataset creation
Video clips for the dataset were sourced from the EPIC-KITCHENS-100 and Kinetics-600 datasets. Using the TagMe platform, annotators formulated questions and answer choices for each category. Each example includes only one correct answer, eliminating ambiguity. Two validation stages were conducted with an annotator overlap of 3, followed by result aggregation. Examples without unanimous annotator agreement underwent additional validation and editing. Post-processing was performed to correct typos. Correct answer options are balanced across classes.
A.3 LabTabVQA
LabTabVQA is a Russian-language question-answering dataset based on images of tables from the medical domain. The dataset includes two types of images: photographs and screenshots (without OCR layers). Each image is paired with a multiple-choice question containing seven answer options, only one of which is correct. The questions are designed to evaluate the capabilities of multimodal LLMs in working with tables presented as images: understanding structure and content, locating and extracting data, analyzing information, etc. All images are anonymized materials from real online consultations on a telemedicine platform.
Motivation
LabTabVQA was created to evaluate the ability of multimodal models to work with tabular information presented in image form, specifically in Russian. Its primary goal is to assess whether such models can understand table structures, interpret their contents, recognize formatting, correlate information, and draw conclusions using only their general knowledge.
The dataset creation and question-generation methodology is not limited to a specific domain and can be extended to include tables from related areas of knowledge. LabTabVQA expands Russian-language benchmarks with a new task category for evaluating models’ ability to analyze tables in terms of content recognition, structural complexity, hierarchy, and data interpretation in end-to-end scenarios.
[Uncaptioned image] Question. What is the sum of the values of all the indicators listed in the heading "Coagulogram"? A. 184.492 B. 169.43 C. 0.92 D. 169.33 E. 184.43 F. 184.44 G. 24.6 Answer. B
Dataset creation
The dataset was built using 697 real images from a telemedicine consultation platform.
Using the GPT-4o Mini model, we annotated images according to two binary criteria:
  • presence of a table in the image;
  • photo or screenshot.
339 images were selected, balanced by image type and table size (also assessed using GPT-4o Mini). For 138 samples, questions were written by experts; for the remaining 201, questions were generated using an AI-agent system composed of the following components:
  1. 1.
    QuestionGenerator (GPT-o4 Mini): generates a candidate question with 7 answer options based on the image and question category;
  2. 2.
    QuestionQualifier (GPT-o4 Mini): identifies the correct answer among the 7 options, or requests regeneration if no correct option is found;
  3. 3.
    Solvers (GPT-4o Mini): at three levels of difficulty (defined by prompts), answer the question and provide reasoning;
  4. 4.
    FeedbackEvaluator (GPT-o4 Mini): analyzes the answers and feedback from the Solvers and decides whether to accept the question or send it back for regeneration (return to step 1).
The generated examples were validated on the TagMe platform (with 3-way overlap) based on the following criteria:
  • the question is based on the table shown in the image;
  • the question does not require domain-specific knowledge (all required information is in the image/table);
  • the question cannot be answered without using the table/image. Similarly, the correct answer was selected by assessors. A correct answer was defined as:
  • the answer proposed by the question generation system, if at least 2 out of 3 assessors agreed with it;
  • the answer chosen by at least 2 out of 3 assessors, even if it differed from the generated answer, provided it was additionally validated by a meta-assessor.
Due to the specifics of the question-generation methodology, the dataset and tasks may be biased toward the GPT-o4 model family.
A.4 RealVideoQA
RealVideoQA is a closed Russian-language question-answering dataset designed for evaluating video-text models (Video-LLMs), comprising questions related to video clips. It comprehensively assesses the following competencies: general video comprehension and detail recognition, possession of common and domain-specific knowledge, the ability to determine the precise order of actions within a video and reconstruct the complete sequence, the capability to count objects and actions over time, as well as the skill to associate actions with their corresponding temporal boundaries in the video. Given a video and a question, the task is to select the single correct answer from four provided options. Correct answers do not require audio track comprehension. All video clips were collected via crowdsourcing and are absent from publicly available sources.
Question. What color is the dome of the tall building in the background on the left? A. Black. B. White. C. Green. D. Blue. Answer. D
Motivation
The majority of published benchmarks in video understanding are focused on English, and currently, no publicly available benchmark exists for the Russian language. The RealVideoQA dataset aims to bridge this gap: it enables the evaluation of how effectively video models can address questions requiring video comprehension (the VideoQA task). This dataset covers the assessment of both basic and advanced model capabilities, including general video comprehension and detail recognition (excluding audio track perception),understanding of diverse question types, and the ability to select the correct answer from provided options.
In the "General Description" category, models must answer questions about the primary action in the video or the foreground object. Questions in the "Attributes and Details" category inquire about specific details or background objects. The "General and Domain Knowledge" category includes questions that necessitate both classical common-sense knowledge and expertise in a specific applied domain (e.g., "In what order should the presented dish be prepared?").The "Action Sequences" category comprises questions testing the understanding of actions in the video, their sequential order, and the ability to reconstruct this sequence. The "Counting" category involves questions assessing the ability to count objects, repetitions of actions over time, and perform basic arithmetic operations with the counts. The "Temporal Intervals" category evaluates the capability to associate actions with specific temporal boundaries (timestamps) within the video. Thus, the dataset tests key competencies essential for the video domain.
Note that the examples do not require audio comprehension, which must be considered during evaluation interpretation.
Dataset creation
Video clips for the dataset were collected via a Telegram bot using crowdsourcing. Annotators formulated questions and answer choices for each category using the TagMe platform. Each example includes only one correct answer, eliminating ambiguity. Two validation stages were conducted with an annotator overlap of 3, followed by result aggregation. Only examples with unanimous annotator agreement were selected. Post-processing was performed to correct typos. Correct answer options are balanced across classes.
A.5 RealVQA
RealVQA is a benchmark for testing the model’s ability to conduct visual question-answering (VQA). The questions are asked in Russian and can relate to a specific object in the image, as well as to the entire image as a whole. The benchmark is built in such a way that it is impossible to answer the question without an image. It is often necessary to conduct logical reasoning in several stages in order to get an answer. A key feature of the dataset is the presence of distractors. Such questions are either about objects that are not present in the image, or there is obviously not enough information to answer the question. The expected behavior of the model in the case of distractor is a message that the question cannot be answered, as well as an indication of the reason why this cannot be done. This is how the model’s resistance to hallucinations is tested.
[Uncaptioned image] Question. Presumably on what day of the week was this photo taken? Answer. on Friday
Motivation
The dataset is designed to evaluate the model’s ability to identify cause-effect relationships and apply logical reasoning based on visual input. The questions are formulated in a way that makes it impossible to answer them without access to the image. Unlike classic VQA datasets that typically assess models’ ability to directly perceive objects (i.e., coarse perception: recognizing simple shapes and colors), this dataset incorporates the most complex types of perception from the AnonymBench taxonomy (understanding relationships between objects and different types of reasoning in particular). A key requirement is that logic or reasoning must be applied to answer the questions. The dataset is intended for state-of-the-art vision and text models that are not only capable of comprehending what is depicted but also performing logical inference. This is a real-world requirement for modern conversational models, as users ask tricky questions about images that have unambiguous answers. Since the questions do not require expert knowledge, the dataset targets everyday scenarios and casual imagery that users might upload in chat applications.
Dataset creation
Image collection was carried out via a Telegram bot under a user agreement ensuring non-disclosure of the photos and user consent. All images were obtained through crowdsourcing, with the condition that the uploaded image must be unique and not previously publicly available online.
The first part of the project involved generating questions–answers pairs using the ABC Elementary platform. The questions were written by AI trainers. These annotators were given an image and tasked with formulating a question and corresponding answer. Emphasis was on complex questions, which were defined as those meeting one of the following criteria: requiring the tracing of causal relationships, understanding or perception of relationships between objects, or requiring additional reasoning to answer. The knowledge required to answer the questions was limited to what is typically covered in the school curriculum and corresponds to general logic, meaning no specialized expertise was necessary.
Additionally, a separate project was created through the ABC Elementary platform for trick questions. The same annotators received photos from the Telegram bot and formulated questions similar to those in the first project, but about objects that were not present in the images.
The third stage of annotation involved verifying the generated questions and answers. Using the ABC Elementary platform, a crowdsourcing approach with an overlap of 3 was employed to validate the created Q&A pairs. The following aspects were checked: 1) the question cannot be answered without the image; 2) the question is neither too general, binary, nor does it require expert knowledge; 3) the answer is unambiguous; 4) the answer adheres to the required format; and 5) the appropriate question type is chosen.
All projects were then aggregated, and the agreed-upon parts were standardized into a unified format. During the verification phase, the question type was further added to the metadata with the following categories: ‘object_properties; logics,other; text_understanding; objects_relationship; knowledge‘. Trick questions comprised 10% of the dataset.
A.6 ruCLEVR
RuCLEVR is a Visual Question Answering (VQA) dataset inspired by the CLEVR Johnson et al. (2017) methodology and adapted for the Russian language.
RuCLEVR consists of automatically generated images of 3D objects, each characterized by attributes such as shape, size, color, and material, arranged within various scenes to form complex visual environments. The dataset includes questions based on these images, organized into specific families such as querying attributes, comparing attributes, existence, counting, and integer comparison. Each question is formulated using predefined templates to ensure consistency and variety. The set was created from scratch to prevent biases. Questions are designed to assess the models’ ability to perform tasks that require accurate visual reasoning by analyzing the attributes and relationships of objects in each scene. Through this structured design, the dataset provides a controlled environment for evaluating the precise reasoning skills of models when presented with visual data.
[Uncaptioned image] Question. Are there any other objects with the same shape as the large metallic object? Answer. no
Motivation
The RuCLEVR dataset was created to evaluate the visual reasoning capabilities of multimodal language models, specifically in the Russian language, where there is a lack of diagnostic datasets for such tasks. It aims to assess models’ abilities to reason about shapes, colors, quantities, and spatial relationships in visual scenes, moving beyond simple language understanding to test compositional reasoning. This is crucial for models that are expected to analyze visual data and perform tasks requiring logical inferences about object interactions. The dataset’s design, which uses structured question families, ensures that the evaluation is comprehensive and unbiased, focusing on the models’ reasoning skills rather than pattern recognition.
Dataset creation
To create RuCLEVR, we used two strategies: 1) generation of the new samples and 2) data augmentation with color replacement. Below, each technique is described in more detail:
Generation of the New Samples: We generated new, unique images and corresponding questions from scratch. This process involved a multi-step process to ensure a controlled and comprehensive evaluation of visual reasoning. First, 3D images were automatically generated using Blender, featuring objects with specific attributes such as shape, size, color, and material. These objects were arranged in diverse configurations to create complex scenes. Questions with the corresponding answers were then generated based on predefined templates, which structured the inquiries into families, such as attribute queries and comparisons. To avoid conjunction errors, we stick to the original format and generate questions in English, further translating them into Russian using Google Translator. After generation, we automatically filtered incorrectly translated questions using the model [^14] https://hf.co/RussianNLP/ruRoBERTa-large-rucola pertained to the linguistic acceptability task. In addition, we checked the dataset for the absence of duplicates.
Data Augmentation with Color Replacement: We also augmented the dataset modifying the images from the validation set of the original CLEVER. Specifically, we developed a script [^15] The link was removed for the review period.to systematically replace colors in questions and images according to predefined rules, thereby creating new augmented samples. This process was initially conducted in English to avoid morphological complexities. Once the questions were augmented, they were translated into Russian and verified for grammatical correctness.
A.7 ruCommonVQA
ruCommonVQA is a publicly available visual question answering dataset in Russian for two types of images: real-world photos and abstract illustrations. The questions are divided into two complexity levels: 1) simple and 2) complex, and categorized by the most frequently occurring types: binary (yes/no), comparative, count-based (how many/much), spatial (where), procedural (how), descriptive (what/which), and subject-based (who). Simple questions can be answered based solely on the visual perception of the image, while complex ones require a step of reasoning. All images in the dataset are standard, sourced from publicly available resources, including real-world or cartoon-style abstract images. ruCommonVQA serves as a foundational VQA dataset for the Russian language and is released under an open and public license.
[Uncaptioned image] Question. Are there any people in the photo? Answer. Yes
Motivation
The dataset addresses the classic foundational Visual Question Answering (VQA) task, similar to English datasets such as VQA Goyal et al. (2017). Currently, there is no publicly available baseline VQA dataset in Russian for evaluating vision-language models. This dataset is designed to assess the core capabilities of models to recognize objects across diverse types of images, understand a variety of question types, and generate answers based on visual input. The question set covers key abilities: understanding objects in the image (Fine-grained Perception, e.g., identification of single instances), overall image perception (Coarse perception), commonsense reasoning, and general knowledge. Images are sourced from public datasets, including COCO Lin et al. (2014) and English-language VQA v2161616https://hf.co/datasets/pingzhili/vqa_v2, which should be considered as limitation when interpreting evaluation results. There is a possibility of indirect data leakage through the images in model training data.
Dataset creation
To construct the dataset, images were sourced from the English-language VQA v2 dataset (which includes data from the COCO Lin et al. (2014) dataset). Using the ABC Elementary platform, annotators created question–answer pairs for the images from scratch. Each image was annotated with 3 questions and with 3-way annotator overlap. The resulting data was then aggregated and filtered both automatically (e.g., removal of overly long answers, typos, formatting issues) and manually. The binary question data was class-balanced.
The second part was created entirely from scratch. To collect images, a Telegram bot was used along with a user agreement that ensured photo confidentiality and confirmed user consent. Images were crowdsourced under the condition that each uploaded image had to be unique and not previously available online or from public sources. In this stage of the project, questions and answers were again generated via the ABC Elementary platform. Questions were written by AI trainers: annotators were provided with an image and instructed to create a question along with a corresponding answer.
A.8 ruEnvAQA
ruEnvAQA is a dataset of multiple-choice and binary-choice questions in Russian. The questions are related to music and non-verbal audio signal understanding. The dataset is based on questions from English-language datasets Clotho-AQA Lipping et al. (2022) and MUSIC-AVQA Li et al. (2022). The questions were translated into Russian and partially modified, while the audio recordings were used in their original form (with length trimming).
The dataset includes 8 types of questions:
  • Original question types from MUSIC-AVQA (approximately half of the questions test expert knowledge about rare instrument sounds, while the rest test general knowledge):
    • ‘Music instrument counting‘: "How many musical instruments are playing in the recording?";
    • ‘Single music instrument detection‘: "Is <instrument_X> playing in the recording?";
    • ‘Double music instrument detection‘: "Is it true that both <instrument_X> and <instrument_Y> are playing in the recording?";
    • ‘Music instrument comparison (louder)‘: "Is it true that <instrument_X> is playing louder than <instrument_Y> in the recording?";
    • ‘Music instrument comparison (longer)‘: "Is it true that <instrument_X> is playing for a longer duration than <instrument_Y> in the recording?";
  • Classes assigned during the editing of CLOTHO-AQA questions (general knowledge questions):
    • ‘Audio scene classification‘ is about understanding the audio scene as a whole, logical inference from multiple details (determining the location or circumstances where the audio was recorded);
    • ‘Audio captioning‘ questions are about understanding specific details of an audio fragment, the order and quantity of events;
    • ‘Sound QA with reasoning‘ questions test audio comprehension with simple reasoning, requiring not only perception of audio signal details but also a step of logical reasoning.
Question. In what location was the recording most likely made? A. at the airport B. at the pier C. at the railway station D. at the bus station Answer. C
Motivation
Compared to the vision-language domain, there are fewer large benchmarks that combine diverse tasks for the evaluation of LALM skills. Examples of such benchmarks include AIR-Bench Yang et al. (2024), AudioBench Wang et al. (2025a), and MMAU Sakshi et al. (2025). Audio understanding tasks can be basically classified into speech analysis, non-verbal signal analysis, and music analysis.
This dataset tests LALMs’ abilities to perceive and analyze non-verbal signals and music by answering questions in Russian about audio recordings of musical compositions and audio scenes from various life situations. The tests include questions of three types:
  • Questions on literal perception of audio events (‘Audio captioning‘ and music questions) test models’ ability to match sequences of events captured in audio, their quantity and duration with their textual description. For example, "How many times did the ball bounce on the floor?" or "Is there a violin playing in the recording?".
  • Questions on audio scene classification (‘Audio scene classification‘) test models’ ability to conduct inductive reasoning, specifically to determine the location and circumstances of audio recording based on event details. For example, if aircraft sounds and announcements are heard in the recording, it was likely made at an airport.
  • Questions with additional reasoning (‘Sound QA with reasoning‘) require additional logical operations with general world knowledge to derive the answer, beyond basic audio information perception. For example, if a cat is meowing in the audio, the question might be: "How do these animals typically move?".
Dataset creation
The dataset is compiled from audio files and questions in equal proportions from two English-language datasets, separately covering the domains of music and non-verbal signals. Questions related to speech understanding are not included in the dataset.
Questions from Clotho-AQA Dataset
The Clotho-AQA Lipping et al. (2022) dataset contains questions about audio with non-verbal signals and minor speech elements, with questions focusing only on non-verbal signals and occasionally on external characteristics of speech, such as volume or speaker gender.
Original questions from the test split were converted to multiple-choice format by generating 3 distractors (incorrect answer options) for each question in addition to the single correct answer from the original dataset. The distractors were generated in English using Llama-3.2-3B-Instruct [^17] https://hf.co/meta-llama/Llama-3.2-3B-Instruct.
Questions, correct answers, and distractors were translated into Russian using DeepL API [^18] https://www.deepl.com/products/api. Questions were translated as a single sequence together with answer options to minimize the impact of synonymy during translation.
The automatically translated questions and answer options, along with corresponding audio files, were reviewed by professional editors (without overlap in annotation) considering the original question formulations. If the original question was unsuitable for translation, the editor posed a new question to the audio, determined the correct answer and distractors. The editor also chose an appropriate question type: Audio scene classification, Audio captioning, or Sound QA with reasoning.
Questions from MUSIC-AVQA
The MUSIC-AVQA Li et al. (2022) dataset consists of video recordings of musical performances and three groups of questions:
  • questions about the audio component of the video, not requiring visual component analysis;
  • questions about the visual content, not requiring understanding of the accompanying audio;
  • questions about audio-visual content, relating simultaneously to both audio and visual parts of the video.
For the ruEnvAQA dataset, only questions related to audio were selected (only test split). The audio component was extracted from each video and used as a standalone wav file.
The selected questions were constructed using templates filled with musical instrument names (22 different instruments):
  • "How many musical instruments are playing in the recording?";
  • "Is <instrument_X> playing in the recording?";
  • "Is it true that both <instrument_X> and <instrument_Y> are playing in the recording?";
  • "Is it true that <instrument_X> is playing louder than <instrument_Y> in the recording?";
  • "Is it true that <instrument_X> is playing for a longer duration than <instrument_Y> in the recording?".
Templates, instrument names, and template answers were translated manually. Questions were selected to balance question types and answers, as well as the musical instruments mentioned in the questions.
The original dataset questions were converted to binary questions. For questions like "How many musical instruments are playing in the recording?", answer options were created as "one" and "several", while other questions were reduced to "yes"/"no" choices. Thus, the resulting dataset has a balance between questions with two and four answer options.
The materials from the original MUSIC-AVQA dataset are protected under the CC BY-NC 4.0191919https://creativecommons.org/licenses/by-nc/4.0/ license, which permits free distribution (including modified materials) for non-commercial purposes.
Question Validation and Audio Processing
Pre-selected questions from both datasets underwent validation by crowdsource annotators with 3-fold overlap. Annotators were presented with an audio, a question, and answer options, and were tasked with selecting all valid answer options to exclude cases with multiple correct answers. Along with validating questions and answers, annotators trimmed the audio to fragments between 5 and 20 seconds in length. If the audio could not be trimmed while maintaining question relevance, the question and audio were excluded.
To obtain aggregated answers, each answer option selection was aggregated using the Dawid-Skene method (each option as an independent variable), after which only questions with a single selected answer option were retained. Subsequently, only annotator answers that matched the aggregated (pseudo-reference) answer were used. The audio fragment in such groups was selected based on the principle of maximum duration, which did not affect the answer since the aggregation grouping was done by question and answer.
A.9 ruHHH-Image
ruHHH-Image is a multimodal dataset designed for Visual Question Answering (VQA) that integrates text and images, with a particular focus on evaluating AI responses through the lens of ethics and safety.
This task checks two key abilities. First, it tests if AI can understand questions with parts from different sources. These sources include both text and images. Second, it evaluates if the AI can choose the best of two answers. The selection is based on ethics or safety categories. The goal is to see if AI can analyze multimodal information. It must then select the most ethical and safe response for users from answer options.
The test is based on two existing datasets. One is the HHH text dataset (1). The other is its Russian version, ruHHH in MERA-text (2). The original categories were Honest, Helpful, and Harmless. ruHHH-Image adds three new ethical categories. These are Empathetic, Etiquette, Open/Impartiality.
Disclaimer. Any videos, questions, and answers in the dataset are not a call to action, condemnation, or insult to anyone or anything. The authors of the dataset pursue exclusively scientific goals, primarily an assessment of the safety and ethical behavior of artificial intelligence models in the proposed situations.
[Uncaptioned image] Question. I have an idea to cheer up my elderly grandma! If I bring her here at night and show her this, do you think shell like it? A. If grandma enjoys horror movies and goes to scary immersive performances on her own, then shell definitely like it. Its totally fine to have such interests at any age. The main thing is to make sure in advance that she actually enjoys this kind of thing. B. This might not be the best way to entertain grandma, but if we cant come up with anything else, it could be worth a try. The main thing is to give her a heads-up - this might not be exactly what someone of her age would enjoy. Answer. A
Motivation
Translated datasets often struggle with different languages and cultures. Ethics is a particularly sensitive area.
ruHHH-Image evaluates models using Russian-language content. This includes texts and photos. It checks if a model can pick the best response. The criteria include honesty, lack of bias, and safety. They also cover empathy, usefulness, and etiquette compliance.
The dataset helps identify problematic responses. These are grouped into the six ethical categories.
In terms of structure, each of the six categories has three subcategories. The dataset balances them equally. There are 33-34 questions per subcategory. This ensures 100 questions per main category.
  1. 1.
    Empathetic Category. Tests formal empathy in three subcategories:
    • animals and plants (inspired by the Voight-Kampff test from Do Androids Dream of Electric Sheep? (1968) by Philip K. Dick),
    • human beings (toward one or a few specific people),
    • society (toward groups or communities).
  2. 2.
    Etiquette Category. Checks adherence to etiquette norms in:
    • place and society (rules for specific locations or groups),
    • time and situations (norms for certain times or scenarios),
    • person (how to behave toward an individual).
  3. 3.
    Harmless Category. Selects the safest answer about situations involving:
    • death,
    • threat (risk of injury or loss),
    • discommode (discomfort, minor inconveniences).
  4. 4.
    Helpful Category. Picks the most useful answer, providing:
    • solutions (direct fixes),
    • prevention (avoiding future problems),
    • development (guidance for growth or benefit)
  5. 5.
    Honest Category. Measures honesty in:
    • truth (factual accuracy),
    • people (avoiding deception),
    • norms (following honesty standards).
  6. 6.
    Open Category. Assesses lack of prejudice toward:
    • groups (based on gender, age, religion, etc.),
    • personal choice,
    • objects, places and actions.
Dataset creation
The dataset was built using images collected through a mobile bot. Annotators checked these images for quality and clarity. Next, questions and answers were created for the images. These covered six ethical categories.
After validation and editing, the categories were split into 18 subcategories. Each main category had three subcategories. This helped capture key aspects of each category. For every image-question pair, annotators provided two to four answer options. They ranked these answers from best to worst. The ranking followed the rules and the requirements of the question’s category.
However, during testing, the model sees only two answers at a time. So some image-question pairs appear up to six times in the dataset. But each time with a different pair of option answers. This method checks if the model ranks answers the same way annotators did.
Limitations
Images and questions reflect Russian-language contexts. Answers align with Russian ethical and cultural views. Not suitable for evaluating global or multicultural ethics. Some sections (Open, Harmless) may go beyond Russian-specific norms into worldwide ones.
A.10 ruHHH-Video
ruHHH-Video is a multimodal dataset that adapts the methodology of ruHHH-Image to the video modality. As the first Russian-specific dataset of its kind, it is designed to evaluate ethical reasoning skills in videos.
A.11 ruMathVQA
ruMathVQA is a multimodal dataset consisting of school math problems presented in the form of images and annotated questions to record the answer in an unambiguous form.
[Uncaptioned image] Annotation. Write the answer as a whole number without specifying units of measurement. Answer. 4
Motivation
The dataset is an open database of tasks for testing the model’s ability to understand pictorial elements from school mathematics and geometry and apply knowledge of school mathematics grades 5-6 and geometry grades 7-9. The peculiarity of this task is to test the models to accurately follow complex mathematical answer formats (annotations), which are fed to the input along with the instruction.
The dataset is intended for SOTA Vision + Text models, which can understand what is depicted and also have some basic knowledge of the school curriculum. The images are presented in the form (the original text of the task is saved inside the picture), which the user can send in the dialog chat to the models in correspondence.
This dataset does not check the course of the solution and does not require deriving reasoning for the problem — the answer to the problem is a short answer with a number/formula. The annotation serves as an instruction for recording an unambiguous short answer to the problem in the form required by the user. Therefore, Accuracy is used as a metric.
Dataset creation
A group of experts with basic knowledge of mathematics was selected for the dataset collection stage. The images for the dataset were drawn by experts — similar to the tasks from school textbooks on mathematics and geometry. The images were drawn in three ways: 1) in an editor on a white sheet using blue or black color; 2) on a white sheet of paper using blue or black color, in uppercase or lowercase letters, with or without the use of drawing tools; 3) on a grid-lined sheet of paper using blue or black color, in uppercase or lowercase letters, with or without the use of drawing tools. The answers to the problems were obtained by solving and discussing each problem by several experts. The annotation, which contains the format for unambiguous recording of the answer to the problem, was manually marked up by an expert by selecting from a list of options for different annotations. A universal question was added to each problem in the instructions: "What is the answer to the problem shown in the picture?"
The dataset obtained in the previous step was validated with overlap by 3 full-time annotators of the ABC Elementary platform. The annotators checked the quality of the images, the answer format and the correctness of the annotation requirements for compliance with the problem question and the answer form. Based on the validation results, if at least one annotator noted an error / poor quality, the data was manually edited.
A.12 ruNaturalScienceVQA
NaturalScienceQA is a multimodal question-answering dataset on natural sciences with basic questions from the school curriculum, based on the English dataset ScienceQA Lu et al. (2022). The dataset includes questions in four disciplines related to natural sciences: physics, biology, chemistry, and natural history. The task requires answering a question based on an image and accompanying context by selecting the correct answer from the options provided. The questions are specifically curated so that it is impossible to determine the correct answer without the image.
Note: A feature of the dataset is that the images used in the tasks may be of relatively low resolution. Thus, the model’s ability to extract information from low-quality images is additionally explored, which is often encountered in applications (e.g., when a user sends a poor-quality screenshot).
[Uncaptioned image] Context. This passage describes a specific growth characteristic in rose plants: Climbing growth and trailing growth are different growth types in roses. Climbing plants have long, bending stems similar to vines. These plants can grow upward, covering fences or walls. Roses with a trailing growth form stay close to the ground, forming low bushes or shrubs. In a group of rose plants, some individuals have a climbing growth habit, while others are trailing. In this group, the gene responsible for growth form has two alleles. The climbing growth allele (G) is dominant over the trailing growth allele (g). This Punnett square shows the cross between two rose plants. Question. What is the expected ratio of offspring with climbing growth to offspring with trailing (bushy) growth? Choose the most likely ratio. A. 4:0 B. 0:4 C. 2:2 D. 3:1 Answer. C
Motivation
The NaturalScienceQA dataset is aimed at evaluating the reasoning abilities of AI models in a multimodal environment. Its goal is to assess models specializing in multimodal reasoning, as the questions involve both textual and visual data and are selected so that they cannot be answered without the image information. It is suitable for models that integrate visual understanding with textual comprehension. The primary users of this dataset are Data Science researchers and developers focused on improving multimodal evaluation, particularly those involved in education, scientific research, and AI-driven tutoring systems. Educators may also find the results valuable in measuring how well AI models can mimic human understanding in educational settings. The questions in the NaturalScienceQA dataset are designed to reflect real-world educational scenarios where students are presented with scientific concepts in visual and textual formats. The dataset evaluates a model’s ability to understand scientific concepts and apply them to solve specific problems. The structure of the questions ensures that models must integrate information from both modalities to determine the correct answer. This design demonstrates that NaturalScienceQA effectively assesses the multimodal reasoning capabilities it aims to test, providing a robust experimental setup for benchmarking AI performance.
Dataset creation
NaturalScienceQA was created based on the English ScienceQA Lu et al. (2022) dataset, a question-answering dataset covering a wide range of scientific disciplines. During the dataset creation process, questions from the test set of the original ScienceQA were selected from four natural science disciplines and manually filtered using the following criteria: 1) the question includes an image and cannot be answered without the accompanying image (relying only on information from the explanatory text), 2) the question is consistent with the Russian educational context and is covered by the school curriculum. Subsequently, the selected questions were translated using the Google Translator API and manually edited to correct errors and inaccuracies from automatic translation. Examples for few-shot learning were obtained similarly but were initially selected from the validation set.
A.13 ruSLUn
RuSLUn (Russian Spoken Language UNderstanding dataset) is a Russian-language dataset for spoken language understanding, designed following the principles of the English SLURP Bastianelli et al. (2020) and the multilingual xSID van der Goot et al. (2021), but with consideration for the cultural and linguistic specifics of Russia. It is intended for evaluating models that map audio recordings directly to semantic representations, including intent detection and slot filling. RuSLUn contains a variety of spoken commands and queries that are typical for Russian users and contexts. The key feature of the dataset is its localization: in addition to being in Russian, it incorporates typical usage scenarios, vocabulary, and contexts, which makes it particularly relevant for developing voice assistants and speech-driven services for Russian-speaking users.
Question.Listen carefully to the audio with the users query, classify the intent the query belongs to, and select all possible slots corresponding to that intent. Words in the slots should have the same morphological form as in the audio, and numbers should be written as text. Answer. {"intent": "RateBook", "slots": { "object_name": "Doctor Zhivago" "rating_value": "three" "best_rating": "six" "rating_unit": "stars" }}
Motivation
Traditionally, the task of spoken language understanding (SLU) is solved in several stages: first, audio recordings are converted into text using automatic speech recognition (ASR), and then the necessary information is extracted from the text using natural language understanding (NLU) technologies. However, this modular approach is susceptible to error accumulation due to ASR inaccuracies, and also requires two separate models or two sequential processing steps, which slows down system performance. The ruSLUn dataset is intended for evaluating audio models capable of directly understanding and interpreting the meaning of audio data in an end-to-end fashion, without an intermediate ASR step. Furthermore, ruSLUn is the first Russian-language dataset in which audio recordings are directly aligned with the corresponding intent and slot annotations. This enables comprehensive research into end-to-end SLU tasks, taking into account the cultural and linguistic specifics of Russian users.
Dataset creation
The dataset was created in two stages: first, text queries were generated and annotated with intents and slots, then, these queries were recorded as audio. The annotation scheme was based on the cross-lingual xSID van der Goot et al. (2021) dataset, which includes 16 intent types and 33 slot types. At the first stage, the validation and test data from xSID were manually translated into Russian by one of the dataset authors. The texts were then adapted to fit the Russian context: locations, names of artists, movies, songs, and restaurants were replaced with popular and recognizable Russian counterparts. These replacements were manually and randomly selected from lists of the most common options. The text data then underwent additional post-processing, including removal of punctuation, conversion of all digits to their word forms, and transforming all text to lowercase. After completing work with the text data, the queries were recorded as audio. Seven speakers of different ages (five women and two men), who were not professional voice actors, were recruited for audio recording. All participants were instructed to record each sentence in a quiet setting, speak in a natural voice, and save each sentence as a separate audio file. Recording took place at home using regular voice recorders, so the audio naturally contains some background noises (such as breaths, shuffling, etc.). The final dataset was manually checked by a moderator to ensure that each audio recording matched the corresponding text data and that the intent and slot annotations were correct.
A.14 ruTiE-Audio
ruTiE-Audio is an emulation of the Turing test in audio format. The dataset consists of a sequence of audio tasks, each accompanied by four possible answers in textual format.
The dataset includes 3 coherent dialogues, each simulating 500 user queries to the model. The model receives an audio input containing tasks and questions, while answer options (4 per task) are provided as text, and the model must choose the most appropriate response. Accordingly, This dataset is designed for evaluating any chat-oriented models capable of processing audio modality.
The tasks test the model’s ability to support a coherent dialogue on naturally changing topics of communication, based on the context of previous interactions.
This dataset is based on the text dataset of the same name from the MERA-text benchmark. In addition to ruTiE-Audio, it is presented in one more version: visual (textual questions about images that are answered in text).
A. Porchbench B. Validol C. Valentina D. Valentine Answer. D
Motivation
This dataset targets models with a large context window (ideally capable of handling dialogue history up to 499 prior turns).
The test has a complex task. The model must not only preserve the context and refer to it in the dialogue, but also have broad linguistic and cultural knowledge: proverbs, nursery rhymes, catchphrases, movie quotes, songs, plays, books, and memes. Moreover, the dataset evaluates spontaneously triggered human-like conversational skills: recognizing irony, the ability to understand and complement a joke, mental arithmetic, spatial reasoning, bilingualism, recognizing and using cause-and-effect relations, avoiding speech traps. Only by using all these skills in a comprehensive manner can one fully "play imitation" according to Turing, that is, adequately participate in a human conversation on equal terms with people.
Please note: during the conversation, the modalities and formats of communication change. The interlocutor can use puns, ask to count the letters in a spoken, not written word, draw your attention to some sound outside the window and wait for your reaction, invite a third person to the conversation, express some opinion or judgment and so on. Therefore, not all prompts are formatted as direct questions. Some are situational utterances or mini audio skits without explicit questions, yet the model must still select the most contextually appropriate response from four given choices. ruTiE-Audio offers 4 answer options for each question.
The test checks the model’s ability:
  • to retain context
  • to support (at the everyday level) a dialogue on any of the main topics (as defined in AnonymBench domains)
  • to recognize and categorize core task types, without which it is impossible to solve the problems of emulating the Turing test (including basic mathematics, ethics, linguistic games, common knowledge, etc.)
  • to navigate in various categories of thinking, including recognizing irony, emotions and intentions of the interlocutor, restoring the essence of the situation based on key elements, etc.
There is also an important limitation for the validity of checking models with the ruTiE-Audio dataset. Since about half of the questions are somehow tied to the immediate context of the emulated "conversation", the next question may suggest the answer to the previous one. So you cannot give the model several tasks from the dialogue at once. Questions are asked strictly one at a time, their order and sequence should not be mixed or changed in any other way.
Dataset creation
The dataset was manually collected by internal experts and then validated. The audio tasks were edited based on scripts written by experts and internal recordings made based on them, previously unpublished online as well. Background noises were sourced from public datasets and custom recordings from the SberDevices studio and various field environments.
A.15 ruTiE-Image
ruTiE-Image is a multimodal emulation of the Turing test, which is an unchangeable sequence of question-answer tasks with the ability to choose an answer. These are 3 coherent dialogues, each dialogue imitates 500 user requests to the model using text and pictures. The model receives answer options (4 for each task) in text form and chooses from them.
The test tasks check the model’s ability to adequately support a dialogue on naturally changing topics of communication, based on the context of previous questions.
The dataset is based on the text dataset of the same name from the MERA-text benchmark. In addition to ruTiE-Image, a similar dataset is presented in one more version: multimodal ruTiE-Audio (questions are submitted to the input in audio format, the model responds with text).
[Uncaptioned image] Question. Hi! Ill call you Ada, and to find out my name, look at the picture and answer who painted them pink - then take the first three letters of that word. So, whats my name? A. Hud B. Sol C. Mal D. Zack Answer. B
Motivation
The dataset is designed to analyze models with a large context window (with a context depth of up to 499 questions).
The test has a complex task. The model must not only preserve the context and refer to it in the dialogue, but also have broad linguistic and cultural knowledge: know proverbs, counting rhymes, catchphrases from films, songs, plays, books, memes. The model must also have skills that are spontaneously actualized in human speech: recognizing irony, the ability to understand and complement a joke, oral arithmetic skills, spatial thinking, bilingualism, recognizing and using cause-and-effect relationships, avoiding speech traps. Only by using all these skills in a comprehensive manner can one fully "play imitation" according to Turing, that is, adequately participate in a human conversation on equal terms with people.
Please note: in a conversation, the modalities of communication often change. The interlocutor can show you a picture, ask you to read the inscription drawn on the wall, refer to a previously shown photo, sometimes invite a third person to the conversation, express some opinion or judgment — and so on. Therefore, the design of a separate task in the ruTiE-Image dialogue is not always designed as a question — it can be designed as a replica-sentence, to which the model needs to choose an adequate reaction. In ruTiE-Image, a task can look like a simple picture sent to the model without an accompanying question — but with suggested reaction options, from which you need to choose the right one. The dataset offers 4 answer options for each question.
The test checks the model’s ability:
  • to retain context,
  • to support (at the everyday level) a dialogue on any of the main subject areas (as defined in AnonymBench domains)
  • to understand the main classes of problems, without which it is impossible to solve the problems of emulating the Turing test (including the simplest mathematics, ethics, linguistic games, general worldview, etc.)
  • to navigate in various categories of thinking, including recognizing irony, emotions and intentions of the interlocutor, restoring the essence of the situation based on key elements, etc.
There is also an important limitation for the validity of checking models with ruTiE. Since about half of the questions are somehow tied to the immediate context of the emulated "conversation", the next question may suggest the answer to the previous one. In this regard, it is not allowed to give the ruTiE model several tasks from the dialogue at once. Questions are asked strictly one at a time, their order and sequence should not be mixed or changed in any other way.
Dataset creation
The dataset was manually collected by internal experts and then verified. The images for the dataset were crowdsourced from previously unpublished mobile photos, ensuring the relevance and modernity of the materials.
A.16 SchoolScienceVQA
SchoolScienceVQA is a Russian-language multimodal dataset inspired by ScienceQA Lu et al. (2022). It evaluates the reasoning capabilities of AI models in a multimodal setting using multiple-choice questions across scientific subjects such as physics, biology, chemistry, economics, history, and earth science. Each question includes an image, text context, and explanation of the correct answer. These components provide a basis for assessing reasoning chains.
[Uncaptioned image] Context. Question. How does the position of the structure highlighted in red in the image change in the mid-latitudes of the oceans in winter compared to summer? A. In winter, it remains at the same depth as in summer B. In winter, it is located deeper than in summer C. In winter, it disappears completely D. In winter, it rises closer to the surface Answer. C
Motivation
SchoolScienceVQA is designed to benchmark AI systems in educational and scientific reasoning tasks requiring both visual and textual understanding. It supports the following use cases:
  • Multimodal Model Evaluation: The dataset requires joint processing of images and text. It is intended for models capable of vision-language reasoning and is unsuitable for unimodal LLMs.
  • Target Audience: Researchers and developers working on multimodal models, especially in the education and tutoring domain. Educators may also use the dataset to measure how well models simulate human-like understanding.
  • Question Content: Questions resemble real-world educational tasks and require true multimodal inference to solve correctly.
Dataset creation
SchoolScienceVQA was developed from scratch based on the methodology of ScienceQA Lu et al. (2022), adapted for Russian cultural and educational context. Domains were adjusted to align with the Russian school curriculum.
Expert annotators from relevant scientific domains created original multimodal examples. Images were produced using original photography, manual illustration, computer graphics, and neural network generation (DALL·E, Stable Diffusion, etc.). All images are novel and not reused from existing datasets. Metadata includes image generation method to support transparency and bias mitigation.
A.17 UniScienceVQA
UniScienceVQA is a multimodal dataset consisting of tasks designed to assess expert knowledge in various fields of science (fundamental, social, and applied sciences, cultural studies, business, health, and medicine). The tasks are presented in the form of images and questions with accompanying annotations. The tasks are divided into three groups based on the response format: 1) short-answer tasks; 2) multiple-choice tasks; and 3) multiple-choice tasks with no correct answer provided.
[Uncaptioned image] Question. What is the order of the automorphism group of the graph shown? Annotation. In your answer, write only the number. Answer. 72
Motivation
The dataset is an open collection of tasks designed to evaluate a model’s ability to understand elements of images from university curricula and professional domains. A distinctive feature of these tasks is testing the model’s capability to provide short and precise answers, as well as to identify the correct answer from multiple-choice options.
The dataset is intended for Vision + Text models that not only understand what is depicted in images but also possess expert knowledge of university-level content.
This dataset does not evaluate the reasoning process or require the model to provide a detailed explanation for solving the task — the answer to the task is a short response in the form of a number or formula. The annotation serves as an instruction for recording an unambiguous short answer to the task in the form required by the user. Therefore, Accuracy is used as the evaluation metric.
Dataset creation
The dataset consists of 25 subdomains, and for data collection in each subdomain, a group of experts with in-depth knowledge in the respective field was involved. The images for the dataset were either drawn or photographed by the experts. The creation of the dataset involved two stages: 1) generating the image, question, and answer; and 2) reviewing the created data. An annotation, which specifies the format for unambiguously recording the answer to the task, was manually added according to the answer. Each task includes a universal instruction: "Read the question and solve the task". As a result, 200-400 tasks were collected for each subdomain.
A.18 WEIRD
WEIRD is an extended version of a binary classification subtask of the original English WHOOPS! Bitton-Guetta et al. (2023) benchmark. The dataset evaluates the ability to detect violations of commonsense. Commonsense violations are situations that contradict the norm of reality Rykov et al. (2025a). For example, penguins can’t fly, children don’t drive cars, guests don’t serve food to waiters, etc. “Weird” and “normal” images are equally distributed in the dataset.
[Uncaptioned image] Question. Is the image strange or normal? A. strange B. normal Answer. A
Motivation
The dataset focuses on evaluating violations of commonsense, and is suitable for the evaluation of any AI models that can analyze images. The main capability that this dataset evaluates is the analysis of visual information and collating it with common sense. Accuracy is the main evaluation metric. Since the dataset evaluates the basic ability to assess plausibility, it will be interesting for any research project as one of the basic stages of the model evaluation pipeline.
Dataset creation
The dataset was created based on the original WHOOPS! Bitton-Guetta et al. (2023), using iterative synthetic generation in the style of Self-Instruct Rykov et al. (2025b). Each sample from the WHOOPS! subset for binary classification is a pair consisting of a “weird” and a “normal” image, along with categories of commonsense violations and image descriptions. To extend the original benchmark, we iteratively generated new categories of commonsense violation and image descriptions using GPT-4o with WHOOPS! samples as a few shots. In addition, we used synthetic descriptions to generate images using DALL-E. Next, we manually filtered out bad images and added good images to the pool. Finally, the pool was used to repeat the generation process and extract new few-shots.
Appendix B Skill taxonomy
This section provides a detailed description of the skill taxonomy for the MERA Multi benchmark, which was introduced in Section 3.2.
Motivation
This taxonomy is designed to cover the skills required from MLLMs to perform common tasks in multimodal domains.
We organize skills into three high-level groups: Perception, Reasoning, and Knowledge.
  • Perception covers extraction of salient information from images/audio/video;
  • Reasoning covers inference over extracted information (often combining multiple cues and background assumptions);
  • Knowledge covers retrieval and application of stored factual or domain information.
This grouping mirrors a standard cognitive decomposition in which perception provides sensory evidence, alignment links that evidence to symbols (e.g., language), reasoning derives new conclusions from evidence and prior beliefs, and knowledge supplies stored factual and conceptual content (long-term memory/semantic representations). In existing multimodal evaluations, similar hierarchical organization appears in benchmark taxonomies (e.g., ConvBench’s perception→reasoning→creativity hierarchy Liu et al. (2024d); MMBench’s fine-grained perception categories Liu et al. (2024e); MMStar’s perception-reasoning-knowledge hierarchy Chen et al. (2024); MME’s perception-cognition Fu et al. (2024)), while knowledge-heavy competence is frequently assessed through QA/exam-style benchmarks such as MMLU/MMMU and cognitively motivated “core knowledge” suites such as CoreCognition Li et al. (2025).
Structure
The core skill set is in the last level of the taxonomy, which is called atomic skills. Each atomic skill has a modality-related version, while particular skills are not applicable to some modalities, e.g. “Image-to-text grounding” is irrelevant for audio modality. In the description that follows, all skill names are provided along with the applicable modality tags ( I A V). To disentangle atomic skill names from aggregation taxonomy categories, the former are marked with ➩ in the following section.
Taxonomy tables (Table 3, Table 7 and Table 8) provide a structured overview of the skill taxonomy along with their coverage by MERA Multi tasks. These tables show the skill hierarchy levels denoted as L1-L5, where L5 are atomic skills (in bold) and L1-L4 are aggregation categories.
Taxonomy Levels Modalities
L1 L2 L3 L4 L5 Image Audio Video
Perception Fine-grained cross-instance perception Overlapping object differentiation Overlapping image differentiation X
Speaker diarization X ruTiE-Audio
AQUARIA
Mutual object localization Spatial object relationship LabTabVQA X CommonVideoQA
ruCommonVQA RealVideoQA
ruCLEVR
ruCLEVR
RealVQA
Temporal object relationship ruTiE-Audio CommonVideoQA
AQUARIA RealVideoQA
Repeating pattern recognition Visual pattern recognition X
Temporal pattern recognition X
Event recognition Object-object interaction ruCommonVQA ruEnvAQA CommonVideoQA
RealVQA AQUARIA RealVideoQA
ruHHH-Image ruHHH-Video
Human-object interaction ruCommonVQA ruTiE-Audio CommonVideoQA
RealVQA AQUARIA RealVideoQA
ruHHH-Image ruHHH-Video
ruTiE-Image
Human-human interaction ruCommonVQA ruTiE-Audio ruHHH-Video
ruHHH-Image AQUARIA
ruTiE-Image
Fine-grained single-instance perception Object recognition Object localization ruCommonVQA X CommonVideoQA
ruCLEVR RealVideoQA
RealVQA
Object recognition ruCommonVQA ruTiE-Audio CommonVideoQA
ruCLEVR ruEnvAQA RealVideoQA
RealVQA AQUARIA ruHHH-Video
ruHHH-Image
ruTiE-Image
Event recognition Object motion recognition ruCommonVQA ruEnvAQA CommonVideoQA
RealVQA AQUARIA RealVideoQA
ruHHH-Image ruHHH-Video
Living things motion recognition ruCommonVQA AQUARIA
RealVQA ruHHH-Video
ruHHH-Image
Pose recognition Non-human pose recognition Animal body pose recognition X
Human pose recognition Human body pose recognition X
Facial expression recognition X
Hand gesture recognition X
Textual grounding Image-to-text grounding Scheme recognition ruMathVQA X
ruNaturalScienceVQA
UniScienceVQA
SchoolScienceVQA
Plot recognition X
Table recognition LabTabVQA X
Text recognition (OCR) LabTabVQA X
ruMathVQA
ruNaturalScienceVQA
UniScienceVQA
SchoolScienceVQA
Audio-to-text grounding Prosody & stress recognition Prosody & stress recognition X
Speech recognition Onomatopoeia X
Speech recognition X ruTiE-Audio
AQUARIA
RuSLUn
Song lyrics recognition X
Media grounding Visual media grounding X
Temporal media grounding X CommonVideoQA
RealVideoQA
Table 7: Skill taxonomy coverage of MERA Multi tasks (perception part). Columns L1-L5 show the skill hierarchy levels, while Image, Audio and Video columns indicate which tasks cover each perception type across different modalities.
B.1 Perception
Perception denotes the set of abilities by which a model extracts task-relevant information from sensory inputs (image, audio, video) and converts it into internally usable representations. In cognitive terms, it is the stage that produces “evidence” from raw signals; downstream competencies (e.g., reasoning or generation) can only be correct if the perceptual evidence is accurate. This dependency is made explicit in hierarchical benchmark taxonomies such as ConvBench Liu et al. (2024d) (perception→reasoning→creativity) and MMStar Chen et al. (2024). In our taxonomy, perception includes both recognition (mapping observable features to concepts, such as objects, events, poses) and coarse localization (relative position without explicit media coordinates).
We define 3 broad categories of perception skills. The dichotomy between single-instance and cross-instance perception has been popularized in MMBench Liu et al. (2024e) and subsequently reused by HR-Bench Wang et al. (2024b). The third category is multimodal alignment which includes skills for matching multimodal representations, including text recognition.
An overview of the perception taxonomy is given in Table 7 and below a detailed skill description is given.
B.1.1 Fine-grained single-instance perception
Perceiving and (optionally) localizing a single salient object/event instance in an input.
Object localization I V
Detecting object positions relative to the image itself (left, right, top, bottom, center) and predicts coarse spatial categories (e.g., left/right/center) without producing media coordinates. In contrast, media grounding predicts locations in the media coordinate system (bounding boxes, segmentation masks, or timestamp spans).
Object recognition I A V
Mapping objective visual and audial features to the concepts in knowledge space, e.g. matching the concept ‘cat’ to meowing sounds and/or visual image of a cat (ears, paws, etc.).
Event recognition (single-instance)
Recognizing actions, events, and states. It involves features that are dynamically perceived (e.g. rolling) or may suggest dynamic characteristics (e.g. color changes), such as those found in motion stop-frames.
Object motion recognition I A V
Living things motion recognition I A V
Recognizing movement types specific to humans and animals.
Pose recognition
The ability to infer articulated body or hand configuration, either as keypoints/joint structure or as gesture/pose classes when the task is categorical. It deserves a separate perceptual skill because many downstream tasks (e.g. action and intention understanding) depend primarily on how a person is posed rather than what objects are present. In practice, this is operationalized by standard pose-estimation datasets such as COCO Keypoints Lin et al. (2014) and MPII Human Pose Andriluka et al. (2014), video benchmarks like PoseTrack Andriluka et al. (2018) for multi-person pose and tracking, and hand-gesture datasets such as HaGRID Kapitanov et al. (2024), for detection- and classification-style gesture recognition.
Body pose recognition I V
Facial expression recognition I V
Hand gesture recognition I V
B.1.2 Fine-grained cross-instance perception
Perceiving and localizing multiple objects or events and the character of their interaction.
Overlapping object differentiation
Separating co-occurring entities whose features interfere.
Overlapping image differentiation I V
Disentangling and separate recognition of visually overlapping objects, .
Speaker diarization A V
Attributing speech to the correct speaker among multiple simultaneous and/or consequent speakers.
Mutual object localization
Perceiving relative relations among multiple entities.
Spatial object relationship I A V
Understanding the relative localization of multiple objects in the scene.
Temporal object relationship A V
Understanding the ordering and temporal relations across events/segments.
Repeating pattern recognition
Perceiving instances that appear as a repeating pattern, e.g. animal flocks, ornaments, rhythmic sounds and moves, also involves perceiving frequency and density. Primary to counting skills.
Visual pattern recognition I V
Patterns present in a single image or video frame, which do not require dynamic perception to acquire.
Temporal pattern recognition A V
Patterns spanning over time and / or perceived only across multiple frames.
Cross-instance event recognition
Perceiving interactions among multiple entities.
Object-object interaction I A V
Human-object interaction I A V
Human-human interaction I A V
B.1.3 Cross-modal alignment
Cross-modal alignment (Grounding) covers abilities that establish correspondences between symbolic descriptions (typically language) and specific elements in a non-text modality, enabling verification (“is the described thing present?”) and reference (“where/when is it?”). In cognitive terms, grounding mediates between perceptual evidence and symbolic reasoning, supporting tasks that require referential linkage rather than mere recognition. We separate two complementary families: (i) Text extraction — recovering textual symbols from media (e.g., OCR in images/video frames; speech recognition and lyrics recognition in audio/video) — and (ii) Media grounding — mapping language to media coordinates (e.g., bounding boxes/masks in images/video, or temporal spans in audio/video). This separation keeps transcription distinct from localization while preserving a unified notion of cross-modal correspondence.
Image-to-text grounding
Involves recognizing text in images, video frames, and more complex visual elements such as tables, charts, and diagrams. Some of the benchmarks targeting these skills together with QA tasks are PlotQA Methani et al. (2020), ChartQA Masry et al. (2022), Misleading ChartQA Chen et al. (2025), ChartBench Xu et al. (2023), and TextVQA Singh et al. (2019).
Text recognition (OCR) I V
Scheme recognition I V
Plot recognition I V
Table recognition I V
Audio-to-text grounding
Similarly to Image-to-text grounding, this skill involves matching sound signals to textual representation, including complex audio elements such as prosody, stress, and lyrics.
Prosody & stress recognition A V
involves recognizing and understanding the rhythm, intonation, and stress patterns in speech.
Speech recognition
Onomatopoeia A V
Onomatopoeia is a language phenomenon where words are used to imitate sounds, such as “buzz” for a bee or “meow” for a cat. Recognizing onomatopoeia is essential for understanding routine language usage. An example of a dataset for onomatopoeia recognition is RWCP-SSD-Onomatopoeia Okamoto et al. (2020).
Speech recognition A V
Song lyrics recognition A V
Media grounding
Grounding to media involves matching textual descriptions to media elements via masking, segmentation, or bounding boxes.
Visual media grounding I V
Operating on bounding boxes, segmentation masks, or pixel coordinates for objects or actions described in natural language. For example, predicting the bounding box of the object in an image or the segmentation mask of the action in a video.
Temporal media grounding A V
Operating on timestamps and media duration. For example, predicting the start and end time of the event in a video or the duration of the action in a video.
B.2 Knowledge
Knowledge denotes a model’s ability to retrieve, apply, and integrate stored factual and conceptual information in multimodal settings. In cognitive science terms, this corresponds to long-term semantic memory (stable representations of entities, relations, norms, and domain concepts) used to interpret perceptual evidence and to support reasoning. In multimodal evaluation, knowledge competence is often probed via QA and exam-style benchmarks (e.g., MMLU/MMMU hendrycks2021measuring; Yue et al. (2024)) that condition questions on images, and via cognitively motivated suites that target foundational concepts from developmental cognition (e.g., CoreCognition Li et al. (2025), which operationalizes “core knowledge” constructs such as object permanence and basic physical/social regularities). Because “knowledge” interacts with both perception and reasoning, we treat it as a distinct group to separate failures of recall/concept access from failures of inference over correctly perceived evidence.
Below we present a detailed description of the knowledge category, which was introduced in Section 3.2.
Knowledge
In MLLM benchmarks, “knowledge” is typically observed as (i) reliance on commonsense/everyday facts in visual contexts, (ii) explicit external/encyclopedic knowledge conditioned on images, and (iii) expert-domain question answering over multimodal inputs.
The differentiation between common and domain knowledge may vary depending on the task and the domain. We formulate the polar points of this distinction as follows: common knowledge is generally agreed upon by the majority of the population and used in everyday life, while expert knowledge is domain-specific and requires specialized knowledge or training.
Common everyday knowledge
Common everyday knowledge I A V
Commonsense and broadly shared facts.
Ethics I A V
Safety/ethics-aligned judgment about harmful/unsafe content or behavior in multimodal prompts. The placement of ethics in the knowledge category and treating ethical norms as stored conventions may be considered a simplification, but it is done to keep minimal the number of major categories.
Domain knowledge
This category may be further subdivided into subject-specific categories as it is done in expert benchmarks, e.g. MMMU Yue et al. (2024).
Common domain knowledge I A V
Subject-specific but broadly accessible knowledge (often school-level or general STEM).
Expert domain knowledge I A V
Specialized, exam- or professional-level knowledge requiring training.
B.3 Reasoning
Reasoning comprises abilities that derive new conclusions from (i) perceptual evidence and (ii) prior knowledge, including uncertain, multi-step, relational, causal, quantitative, and counterfactual inference. In cognitive terms, reasoning corresponds to controlled inference processes that operate on representations supplied by perception and memory, producing answers that are not directly observable in the input. We include both classical forms (deductive/abductive/inductive inference) and task-driven “holistic” judgments (e.g., topic, style, provenance, media characteristics) when they require integrating multiple cues across an input rather than naming a single localized entity. This placement is consistent with benchmark taxonomies that separate fine-grained recognition from higher-level judgments—for example, MMBench distinguishes coarse perception categories such as image topic/quality from fine-grained object-level recognition Liu et al. (2024e). Analogously, audio evaluation benchmarks emphasize media-level characterization and instruction-guided understanding (AudioBench Wang et al. (2025a), AIR-Bench Yang et al. (2024), MMAU Sakshi et al. (2025)).
Taxonomy Level Modality
L1 L2 L3 L4 L5 Image Audio Video
Reasoning Inductive reasoning Attribute recognition Coarse attribute recognition Generated content detection
Source characterization
Media characteristic understanding
Speech emotion recognition AQUARIA
Music emotion recognition AQUARIA
Melodic structure interpretation
Topic understanding ruTiE-Image ruTiE-Audio
Style & genre understanding RealVQA AQUARIA
Scene understanding ruCommonVQA ruTiE-Audio CommonVideoQA
RealVQA ruEnvAQA RealVideoQA
ruHHH-Image AQUARIA ruHHH-Video
ruTiE-Image
Object attribute recognition Physical property understanding ruCommonVQA ruEnvAQA CommonVideoQA
ruCLEVR AQUARIA RealVideoQA
RealVQA
UniScienceVQA
SchoolScienceVQA
WEIRD
Object function understanding ruCommonVQA ruEnvAQA CommonVideoQA
RealVQA RealVideoQA
ruHHH-Image ruHHH-Video
WEIRD
Identity & emotion understanding ruCommonVQA AQUARIA
WEIRD
Other inductive reasoning Other inductive reasoning
Deductive reasoning Weirdness understanding WEIRD
Analogical reasoning ruTiE-Image ruTiE-Audio
Other deductive reasoning
Abductive reasoning Hypothetical reasoning RealVQA AQUARIA CommonVideoQA
Cause & effect understanding RealVQA AQUARIA CommonVideoQA
RealVideoQA
Quantitative reasoning Counting Static counting LabTabVQA X CommonVideoQA
ruCommonVQA RealVideoQA
ruCLEVR
ruNaturalScienceVQA
RealVQA
UniScienceVQA
SchoolScienceVQA
ruTiE-Image
Temporal counting X ruTiE-Audio CommonVideoQA
ruEnvAQA RealVideoQA
AQUARIA
Mathematical reasoning Mathematical reasoning ruMathVQA ruTiE-Audio CommonVideoQA
ruNaturalScienceVQA RealVideoQA
RealVQA
UniScienceVQA
SchoolScienceVQA
ruTiE-Image
Other reasoning Critical thinking
Counterfactual robustness RealVQA
Problem decomposition ruMathVQA
ruNaturalScienceVQA
RealVQA
UniScienceVQA
SchoolScienceVQA
Comparative reasoning RealVQA ruEnvAQA
UniScienceVQA AQUARIA
SchoolScienceVQA
Table 8: Skill taxonomy coverage of MERA Multi tasks (reasoning part). Columns L1-L5 show the skill hierarchy levels, while Image, Audio and Video columns indicate which tasks cover each reasoning type across different modalities.
An overview of the reasoning taxonomy is given in Table 8 and below a detailed skill description is provided.
B.3.1 Inductive reasoning
Attribute recognition
The distinction between coarse attribute recognition and object attribute recognition is inspired by MMBench’s Liu et al. (2024e) categorization into coarse vs. fine-grained perception. We put coarse perception into reasoning category as most of the related tasks require inference from the perceived information about the whole sample.
Coarse attribute recognition
Global or “holistic” characterization of a sample (quality, topic, provenance, modality-level traits), without requiring fine object-level localization or multi-entity comparison.
Topic understanding I A V
Style & genre understanding I A V
Scene understanding I A V
Generated content detection I A V
Detection of whether content is synthetic vs real.
Media characteristic understanding I A V
Recognize broad media-level traits (quality, augmentations). Characterizing/attributing the origin or type of synthesis (e.g., illustration type, camera quality, sound interference).
In vision-language evaluation, MMBench Liu et al. (2024e) explicitly includes Image Quality and Image Topic as coarse perception categories. In audio(-text) evaluation, media-level characterization is a core target of AudioBench Wang et al. (2025a) (speech understanding, audio scene understanding, and paralinguistic “voice understanding”), AIR-Bench Yang et al. (2024) (speech/natural sounds/music comprehension under instruction following), and MMAU Sakshi et al. (2025) (multi-task audio understanding and reasoning across speech/environment/music).
Speech emotion recognition A V
Music emotion recognition I A V
Melodic structure interpretation I A V
Object attribute recognition
Attribute-centric understanding grounded in entities (objects, people, instruments), including physical properties, functions, and identity-related cues.
Physical property understanding I A V
Infer physical properties (e.g., material/shape/size/texture; or physical dynamics in video).
Object function understanding I A V
Infer what an object is for and how it is used (affordances, intent).
Identity & emotion understanding I A V
Recognize identity-related cues (who) and affective state (facial expression, vocal affect).
B.3.2 Deductive reasoning
Rule- or structure-driven inference where conclusions follow from constraints (logical, relational, symbolic, or structural).
Weirdness understanding I A V
Detect incongruity, humor, anomalies, or “oddness” that violates expectations.
Analogical reasoning I A V
Solve problems by mapping relational structure between situations.
Other deductive reasoning I A V
General logical/relational deduction (multi-hop, constraints).
B.3.3 Abductive reasoning
Inference to the best explanation: forming hypotheses and causal attributions consistent with observations.
Hypothetical reasoning I A V
Reason about “what if” interventions and unseen outcomes.
Cause & effect understanding I A V
Identify causal responsibility and predict effects of interactions over time.
B.3.4 Quantitative reasoning
Counting
Static counting I A V
Count entities in a single frame/image. Complementary to pattern recognition skills that focus on repeated structures.
Temporal counting A V
Count events and objects over time in audio/video streams. Complementary to temporal pattern recognition skills that focus on repeated structures.
Mathematical reasoning
Solve math problems grounded in visual contexts (diagrams, plots, tables, scenes) or multimodal inputs. MathVista Lu et al. (2024) is explicitly proposed to evaluate mathematical reasoning of foundation models in visual contexts.
Mathematical reasoning I A V
B.3.5 Other reasoning
Critical thinking I A V
Evaluate evidence, reconcile inconsistencies, and make defensible judgments under ambiguous multimodal context.
Counterfactual robustness I A V
Maintain correct reasoning when interventions are applied (e.g. answer option in MCQ do not contain the correct answer, or the question relates to an entity not present in the image).
Problem decomposition I A V
Solve by breaking into intermediate steps; multi-step chains across modalities. ScienceQA Lu et al. (2022) is explicitly proposed to evaluate problem decomposition skills of foundation models.
Comparative reasoning I A V
Answer questions requiring comparisons (more/less, same/different) across instances or across time.
Appendix C Data leakage details
Goal. Given a multimodal example $x=(t,m)$ where $m$ is the paired modality (image/video/audio), $t$ is the text, estimate the probability that a target model was trained on $x$.
C.1 Setup.
We begin by creating a controlled environment to simulate data leakage. For a given modality, we take a base model, $Model$. We then create a leaked version, $Model_{leak}$, by fine-tuning the base model on a subset of our benchmark data using SFT-LoRA for the selected modality.
C.2 Neighbor Generation and Feature Extraction.
For each original data point $(t,m)$ we generate $K=24$ perturbed "neighbors". We apply four distinct perturbation techniques (masking and predicting the masks with Fred-T5 model [^20] https://hf.co/ai-forever/FRED-T5-1.7B, Zmitrovich et al. (2024), deletion, duplication, and swapping of random words) to the text $t$ with each technique applied 6 times. The modality data $m$ remains unchanged.
For each original text $t$ and its neighbors $t_{k}^{\prime}$ we extract their text embeddings using a fixed encoder:
$e=E(t),\quad e_{k}^{\prime}=E(t_{k}^{\prime})$
where $E$ is intfloat/e5-mistral-7b-instruct [^21] https://hf.co/intfloat/e5-mistral-7b-instruct. It used to be SoTA on MTEB benchmark Muennighoff et al. (2022) model.
Subsequently, we compute the multimodal loss for both models $Model$ and $Model_{leak}$ on both the original and neighbor data points:
$\mathcal{L}=L(Model,t,m),\quad\mathcal{L}_{k}^{\prime}=L(Model,t_{k}^{\prime},m)$
C.3 Detector Training
The core of MSMIA is a binary classifier trained to distinguish between models that have and have not seen the data. For each neighbor $k$ we create two training examples by computing the feature differences:
$\Delta\mathcal{L}=\mathcal{L}-\mathcal{L}_{k}^{\prime},\quad\Delta e=e-e_{k}^{\prime}$
These feature vectors are paired with labels $y\in\{0,1\}$ indicating whether the losses came from ModelModel ($y=0$) or $Model_{leak}$ ($y=1$). This process yields 48 training triplets $(\Delta\mathcal{L},\Delta e,y)$ per original data point. The MSMIA detector, $f_{MSMIA}$ is trained to predict the probability $p=f_{MSMIA}(\Delta\mathcal{L},\Delta e)$ that the input features originate from a model that has been trained on the target data.
C.4 Inference and Evaluation
To infer if a target $TestModel$ has been trained on a specific data point $(t,m)$, we repeat Step 2 to compute the loss and embedding differences for this model. We then compute the leakage score $S$ for the data point by taking the average probability output by the detector over all $K$ neighbors:
$S(t,m)=\frac{1}{K}\sum_{k=1}^{K}f_{MSMIA}(\Delta\mathcal{L}_{k},\Delta e_{k})$
To get the probability estimation for the entire dataset, the $S$ scores are averaged over all dataset samples.
Origin Model Test Model AUC-ROC
Qwen2.5-VL-3B-Instruct Qwen2.5-VL-3B-Instruct 96.2
Qwen2.5-VL-3B-Instruct Qwen2-VL-7B-Instruct 86.0
Qwen2.5-VL-3B-Instruct Qwen2.5-VL-7B-Instruct 88.0
Qwen2.5-VL-3B-Instruct llama3-llava-next-8b-hf 90.2
Qwen2.5-VL-3B-Instruct gemma-3-4b-it 65.8
Qwen2.5-VL-3B-Instruct gemma-3-12b-it 67.9
Qwen2-VL-7B-Instruct Qwen2.5-VL-3B-Instruct 78.0
Qwen2-VL-7B-Instruct Qwen2-VL-7B-Instruct 96.2
Qwen2-VL-7B-Instruct Qwen2.5-VL-7B-Instruct 80.5
Qwen2-VL-7B-Instruct llama3-llava-next-8b-hf 78.0
Qwen2-VL-7B-Instruct gemma-3-4b-it 77.7
Qwen2-VL-7B-Instruct gemma-3-12b-it 73.7
Qwen2.5-VL-7B-Instruct Qwen2.5-VL-3B-Instruct 92.8
Qwen2.5-VL-7B-Instruct Qwen2-VL-7B-Instruct 93.1
Qwen2.5-VL-7B-Instruct Qwen2.5-VL-7B-Instruct 98.1
Qwen2.5-VL-7B-Instruct llama3-llava-next-8b-hf 95.8
Qwen2.5-VL-7B-Instruct gemma-3-4b-it 95.4
Qwen2.5-VL-7B-Instruct gemma-3-12b-it 94.5
llama3-llava-next-8b-hf Qwen2.5-VL-3B-Instruct 94.6
llama3-llava-next-8b-hf Qwen2-VL-7B-Instruct 90.0
llama3-llava-next-8b-hf Qwen2.5-VL-7B-Instruct 96.6
llama3-llava-next-8b-hf llama3-llava-next-8b-hf 97.7
llama3-llava-next-8b-hf gemma-3-4b-it 99.1
llama3-llava-next-8b-hf gemma-3-12b-it 99.5
gemma-3-4b-it Qwen2.5-VL-3B-Instruct 76.0
gemma-3-4b-it Qwen2-VL-7B-Instruct 71.5
gemma-3-4b-it Qwen2.5-VL-7B-Instruct 85.2
gemma-3-4b-it llama3-llava-next-8b-hf 86.5
gemma-3-4b-it gemma-3-4b-it 99.4
gemma-3-4b-it gemma-3-12b-it 98.7
gemma-3-12b-it Qwen2.5-VL-3B-Instruct 84.1
gemma-3-12b-it Qwen2-VL-7B-Instruct 81.3
gemma-3-12b-it Qwen2.5-VL-7B-Instruct 91.2
gemma-3-12b-it llama3-llava-next-8b-hf 93.3
gemma-3-12b-it gemma-3-4b-it 99.4
gemma-3-12b-it gemma-3-12b-it 99.7
Table 9: AUC-ROC MSMIA performance metrics for various evaluated Image MLLMs.
Origin Model Test Model AUC-ROC
Qwen2.5-VL-3B-Instruct Qwen2.5-VL-3B-Instruct 95.9
Qwen2.5-VL-3B-Instruct Qwen2.5-VL-7B-Instruct 99.5
Qwen2.5-VL-3B-Instruct LLaVA-NeXT-Video 91.7
Qwen2.5-VL-3B-Instruct LLaVA-NeXT-Video-DPO 91.2
Qwen2.5-VL-7B-Instruct Qwen2.5-VL-3B-Instruct 98.7
Qwen2.5-VL-7B-Instruct Qwen2.5-VL-7B-Instruct 100.0
Qwen2.5-VL-7B-Instruct LLaVA-NeXT-Video 96.5
Qwen2.5-VL-7B-Instruct LLaVA-NeXT-Video-DPO 95.7
LLaVA-NeXT-Video Qwen2.5-VL-3B-Instruct 63.7
LLaVA-NeXT-Video Qwen2.5-VL-7B-Instruct 71.5
LLaVA-NeXT-Video LLaVA-NeXT-Video 100.0
LLaVA-NeXT-Video LLaVA-NeXT-Video-DPO 100.0
LLaVA-NeXT-Video-DPO Qwen2.5-VL-3B-Instruct 53.6
LLaVA-NeXT-Video-DPO Qwen2.5-VL-7B-Instruct 56.2
LLaVA-NeXT-Video-DPO LLaVA-NeXT-Video 100.0
LLaVA-NeXT-Video-DPO LLaVA-NeXT-Video-DPO 100.0
Table 10: AUC-ROC MSMIA performance metrics for various evaluated Video MLLMs.
Origin Model Test Model AUC-ROC
Qwen2-Audio-7B-Instruct Qwen2-Audio-7B-Instruct 87.7
Qwen2-Audio-7B-Instruct Qwen-Audio-Chat 76.0
Qwen-Audio-Chat Qwen2-Audio-7B-Instruct 61.3
Qwen-Audio-Chat Qwen-Audio-Chat 100.0
Table 11: AUC-ROC MSMIA performance metrics for various evaluated Audio MLLMs.
We report AUC-ROC for binary classification (leaked vs. clean) as shown in Tables 9, 11, 10. There the Origin Model is the model used to train MSMIA. Test Model is the model whose losses are used to test MSMIA (predict whether the data sample was used to train Test Model or not).
Appendix D LLM-as-a-judge details
D.1 Data collection
Our dataset comprises (question, gold answer, model prediction) triplets sourced from Russian-language benchmarks, including MERA Multi, with model sizes spanning 2B to 110B parameters. Human annotators label the semantic correctness of each prediction, strictly ignoring surface form. To ensure label quality, only items with 100% inter-annotator agreement are used for model training and testing.
Total Open Generation Multiple Choice
Train set 62,580 22,538 40,042
Test set 8,570 5,819 2,751
Table 12: Overview of the datasets used for training and evaluating the judge model. The columns indicate the total number of examples and their distribution across two task formats: Open Generation (free-form model response) and Multiple Choice (discrete answer selection)
We apply synthetic augmentations (e.g., refusals, repetitions) to improve robustness. A critical measure to prevent bias was splitting the data by source dataset, ensuring no dataset appears in both train and test splits. The final dataset composition and class balance are detailed in Table 12.
D.2 Training models
The tasks are formulated as a binary classification problems. Inputs are linearized into the format question [SEP] gold answer [SEP] model prediction and packed into the maximum models’ contexts. We train encoder-based models, adding a linear classification head on top of the pre-trained backbone. The models are fully fine-tuned and optimized using a cross-entropy loss, with class weights applied to mitigate potential dataset imbalance.
Parameter Value
Learning Rate 2e-5
Batch Size 32
Number of Epochs 3
Weight Decay 0.01
Optimizer AdamW
LR Scheduler Linear
Max Sequence Length 512
Precision BF16
Early Stopping Metric F1 Score
Table 13: Training configuration summary for embedding-based classification
Optimization is performed with mixed precision and early stopping based on the F1 score on the development set, constrained to a single A100 80GB GPU. A summary of the final training configuration is provided in Table 13.
D.3 Model selection
We compare fine-tuned encoder-based models against zero-shot decoder baselines to identify the optimal judge architecture. The decoder models are prompted to output either “0” or “1” without any task-specific fine-tuning. While this approach is flexible, its reliability is limited by the fact that decoder generations are unconstrained; they may not always produce a valid classification, making score interpretation ambiguous.
In our experiments, encoder models proved to be more suitable for this binary classification task. Their architectural design naturally provides probability distributions over the two classes (correct/incorrect), ensuring deterministic scoring. Furthermore, they offer significant practical advantages, being generally smaller, faster on a single GPU, and even capable of handling long contexts (useful for judging the answers with reasoning or chain-of-thought elements). A full model comparison with bootstrap statistics is provided in Table 14.
Model Parameters Context Length Samples/sec F1 Recall Precision Pearson EM rate
Encoders RuModernBERT-base 150M 8,192 1800 0.964 $\pm$ 0.002 0.975 0.955 0.940 0.997
RuModernBERT-small 35M 8,192 2000 0.944 $\pm$ 0.003 0.962 0.927 0.879 0.982
Qwen3-Embedding-0.6B 600M 32,768 1850 0.951 $\pm$ 0.003 0.985 0.920 0.936 1.000
FRIDA 823M 512 70* 0.838 $\pm$ 0.005 0.913 0.774 0.765 0.925
embeddinggemma-300m 303M 2,048 250* 0.915 $\pm$ 0.004 0.965 0.871 0.825 0.997
Giga-Embeddings-instruct 3.45B 4,096 40* 0.966 $\pm$ 0.002 0.984 0.948 0.946 1.000
Decoders pollux-judge-7b 7.61B 131072 20 0.823 $\pm$ 0.005 0.788 0.860 0.732 0.873
T-lite-it-1.0 7.61B 131072 52 0.844 $\pm$ 0.005 0.968 0.747 0.755 0.993
Qwen3-0.6B 752M 32768 442 0.030 $\pm$ 0.003 0.016 0.323 -0.010 0.004
Qwen3-1.7B 2.03B 32768 205 0.590 $\pm$ 0.006 0.990 0.420 0.300 0.997
gpt-oss-20b 21.5B 131072 28 0.939 $\pm$ 0.003 0.968 0.912 0.905 1.000
Table 14: Summary results for judge models on the test set. F1, Recall and Precision report binary classification quality; Pearson is the Pearson correlation between binary predictions and ground-truth labels; EM rate is the share of class-1 predictions on the subset of examples with an exact string match between the prediction and the gold answer. An asterisk (*) next to throughput indicates the model could not be served via vLLM and the speed was measured with the HuggingFace Trainer instead
D.4 Model deployment
The selected judge model is deployed using vLLM for high-throughput inference on a single A100 40GB GPU. This provides optimal batch processing speed for rapid benchmark evaluation while maintaining quality. Our codebase implements a score(q, ref, pred) → {0,1} API that handles individual scoring and aggregation. We also publicly provide the trained model with weights on HuggingFace Hub  [^22] https://huggingface.co/MERA-evaluation/MERA_Answer_judge.
D.5 Model analysis
To validate our LLM-as-a-judge, we performed a sanity check demonstrating high heuristic consistency: the model aligns with human labels in 99.6% of cases where Exact Match (EM) is 1, and 97.7% when EM is 0 but humans deem the response correct. This reliability extends across task types, with agreement rates of 98.7% for multiple-choice and 96% for free-form generation. Statistical analysis confirms the judge is resilient to common biases, showing near-zero Pearson and Spearman correlations with answer length ($r\approx 0.001$, $\rho\approx-0.008$) and gold label position ($r/\rho\approx-0.0005$). Qualitative analysis reveals that errors primarily occur in tasks involving complex LaTeX formatting or multi-step reasoning. Since our model is designed as a binary classifier without a rationale-generation component, it occasionally struggles with the ambiguous justifications or symbolic intricacies inherent in these specialized domains.
Appendix E Block prompts analysis
E.1 Prompts formulations
There are 13 blocks:
  • Attention hook (greeting / draw attention).
  • General task description (task specific).
  • Input data description (enumerate modalities).
  • Action on data (e.g., “Solve the task using the images…”).
  • Optional task specifics (helpful but nonessential context).
  • Textual question.
  • Answer options (if multiple choice).
  • Call to solve (explicit request to answer).
  • Reasoning request (ask for thinking before final answer; present in 5 prompts).
  • Reasoning format (how to present the reasoning).
  • Answer format (how to present the final answer).
  • Time limitation (e.g., “you have 10 minutes”).
  • Final call to action (e.g., “get started”).
The blocks are combined with the Python script, following the task prompts configuration file. This file contains an explicit description of all ten prompts. Example of the prompt configuration:
# prompt_config.yaml
prompt_4:
attention_hook: "informal_request"
task_description: "informal_request"
input_data: "default"
processing_data: "informal_request"
context_intro: "in_dataset"
task_context: "default"
question: "default"
answer_options: "default"
solution_motivation: "informal_request"
reasoning_motivation: "none"
reasoning_format: "none"
answer_format: "informal_request"
limitations: "informal_request"
answer_motivation: "informal_request"
The block formulations are fixed. The only variable blocks are task_description and task_context that are task-specific. All prompt formulations are imputed in the dataset on HuggingFace Hub (key “instruction” of each data sample). The blocks distribution is fixed: 5 prompts with reasoning_motivation and reasoning_format blocks to enable answer rationale, 1 prompt with no answer_format (“zero prompt” that provides the minimal version of the task - only placeholders for multimodal data, the question and answer options if any).
The example of the prompt with all blocks are as follows.
“Zero prompt”:
Image: <image> Question: {question}
Prompt with specific answer format (RuSLUn):
The dataset for the task includes the following prompt: Not all slots are necessarily present in the request. Audio file: <audio> Question: {question} Please solve the task based on the above and briefly formulate your answer. {annotation}
Example of the prompt with “reasoning” blocks:
The dataset for the task includes the following prompt: The question is directly related to the content of the image and requires not only recognition of individual elements, but also understanding of the relationships between the elements (objects) in the image. If there is insufficient information to answer the question, for example, if the object in question is missing from the image, then you must honestly answer that the question cannot be answered and indicate the reason. Image: <image> Question: {question} Please solve the task based on the above and briefly formulate your answer. Please think about the solution and describe your thought process in detail. Write your reasoning after the word REASONING, briefly explaining how you arrived at your final answer. Please provide a brief answer to the question. Please do not write anything else, do not elaborate, do not engage in dialogue, and do not explain your answer. Please write your final answer after the word ANSWER.
E.2 Statistical analysis
To analyze the effect that a prompt formulation has on the metrics, we fit OLS with the following specification:
$metric\sim C(prompt)+C(model)+C(engine)$ (5)
In fact, it is essentially the same as:
$\begin{split}metric_{i}=\alpha+\sum_{p=2}^{10}\beta_{p}\{prompt=p\}+\\ \sum_{m}\gamma_{m}\{model=m\}+\sum_{e}\delta_{e}\{engine=e\}\end{split}$ (6)
Where:
  • Prompt is a specific prompt formulation (one of 10).
  • Model is the model name - the model used to infer the data with the prompt and the metric metric.
  • Engine is the inference backend used for evaluation (one of transformers (Wolf et al., 2020) or vllm [^23] https://github.com/vllm-project/vllm).
  • Frames. Additional categorical variable used only for video modality evaluations - the video is uniformly split into $N$ frames and the vision LLM used to make evaluation on the dataset.
  • Domain. Additional categorical variable used only for UniScienceVQA, SchoolScienceVQA, ruCommonVQA. These datasets are split into separate domains (subsets). This split may affect the metrics (one domain may be “harder” than another).
For each prompt $p$, we test $H_{0}:\beta_{p}=0$ against $H_{1}:\beta_{p}\neq 0$. We mark a prompt’s effect as statistically significant when the two-sided p-value <0.05<0.05.
Figure 2: The relative (with regard to baseline prompt (0)) effects of different formulations of prompts for each dataset. There are ten different formulations of prompts for one dataset, hence nine corresponding bars (one formulation is baseline category). Red bars reflect statistically significant (at 95% confidence level) effects.
Figure 2 demonstrates the results of statistical analysis of the prompts formultations effects on the Judge Score metric. There are three datasets that have been omitted: RuSLUn implies structured output with rather strict metric that tends to show zero score, ruTiE-Audio and ruTiE-Image datasets scores are too sensible to prompt formulation due to the datasets design [^24] Both datasets consist of three sequential dialogues of 500 questions in a row. The answer for the question $X$ lays in the text of the questions >X>X. This way we cannot reliably separate this effect from the pure effect of the prompt formulation which may lead to incorrect analysis and conclusions.
Takeaways:
  • No single prompt dominates. Different datasets favor different formulations; the prompt that helps in one case can hurt in another. This undermines any one-prompt-fits-all strategy.
  • Magnitude varies widely across datasets. Some tasks show large shifts (on the order of 0.1–0.2 in the metric), while others exhibit near-zero effects (often still significant when variance is low). Prompt sensitivity is therefore task-dependent.
  • Reasoning format does not mean lower scores. Prompts that explicitly encourage model to provide chain-of-though rationale for the final answer do not always get lower scores. First, this means that the LLM-as-judge model is capable of coping with reasoning format. Second, some models breach the answer format prescribed in the task instruction.
  • Single prompt leads to bias. Only four tasks (notable, three of them are from video modality) demonstrate no red bars - no statistically significant prompt formulations compared to the baseline one. The other 12 datasets tend to be prompt sensitive, so the design choice to distribute prompts uniformly and interpret dataset metrics as an average over formulations is empirically supported.
Appendix F Baselines Details
F.1 Model Baselines Details
In this section, the list of baseline models is provided. Tables 15, 16, 17 represent models for image, audio, and video modalities respectively.
Model Parameters Context length Hugging Face Hub link Citation
GPT 4.1 N/A 1000K GPT 4.1
Phi-3.5-vision-instruct 4B 128K microsoft/Phi-3.5-vision-instruct Abdin et al. (2024)
Phi-4-multimodal-instruct 6B 128K microsoft/Phi-4-multimodal-instruct Microsoft et al. (2025)
Qwen2.5-Omni-3B 6B 32K Qwen/Qwen2.5-Omni-3B Xu et al. (2025)
Qwen2.5-Omni-7B 11B 32K Qwen/Qwen2.5-Omni-7B
Qwen2-VL-2B-Instruct 2B 32K Qwen/Qwen2-VL-2B-Instruct Wang et al. (2024a)
Qwen2-VL-7B-Instruct 7B 32K Qwen/Qwen2-VL-7B-Instruct
Qwen2-VL-72B-Instruct 72B 32K Qwen/Qwen2-VL-72B-Instruct
Qwen3-VL-2B-Instruct 2B 262K Qwen/Qwen3-VL-2B-Instruct Yang et al. (2025)
Qwen3-VL-8B-Instruct 9B 262K Qwen/Qwen3-VL-8B-Instruct
Qwen2.5-VL-3B-Instruct 3B 128K Qwen/Qwen2.5-VL-3B-Instruct Bai et al. (2025)
Qwen2.5-VL-7B-Instruct 7B 128K Qwen/Qwen2.5-VL-7B-Instruct
Qwen2.5-VL-32B-Instruct 32B 128K Qwen/Qwen2.5-VL-32B-Instruct
Qwen2.5-VL-72B-Instruct 72B 128K Qwen/Qwen2.5-VL-72B-Instruct
gemma-3-12b-it 12B 128K google/gemma-3-12b-it Team et al. (2025a)
llava-1.5-13b-hf 13B 4K llava-hf/llava-1.5-13b-hf Liu et al. (2024b)
llava-next-72b-hf 72B 8K llava-hf/llava-next-72b-hf Li et al. (2024)
llava-next-110b-hf 110B 32K llava-hf/llava-next-110b-hf
InternVL3-9B 9B 32K OpenGVLab/InternVL3-9B Zhu et al. (2025)
granite-vision-3.3-2b 3B 128K ibm-granite/granite-vision-3.3-2b Team et al. (2025b)
SmolVLM-Instruct 2B 16K HuggingFaceTB/SmolVLM-Instruct Marafioti et al. (2025)
MiniCPM-o-2_6 9B 32K openbmb/MiniCPM-o-2_6 Yao et al. (2024)
Table 15: General information about vision (image) modality baseline models.
Model Parameters Context length Hugging Face Hub link Citation
Qwen3-Omni-30B-A3B-Instruct 35B 8K Qwen/Qwen3-Omni-30B-A3B-Instruct
Qwen2.5-Omni-3B 6B 32K Qwen/Qwen2.5-Omni-3B Xu et al. (2025)
Qwen2.5-Omni-7B 11B 32K Qwen/Qwen2.5-Omni-7B
Qwen2-Audio-7B-Instruct 7B 32K Qwen/Qwen2-Audio-7B-Instruct Chu et al. (2024)
Qwen2-Audio-7B 8B 32K Qwen/Qwen2-Audio-7B
audio-flamingo-3-hf 9B 32K nvidia/audio-flamingo-3-hf Goel et al. (2025)
SeaLLMs-Audio-7B 8B 8K SeaLLMs/SeaLLMs-Audio-7B Zhang et al. (2024)
ultravox-v0_2 8B 8K fixie-ai/ultravox-v0_2 Team (2024)
ultravox-v0_3 8B 8K fixie-ai/ultravox-v0_3
ultravox-v0_3-llama-3_2-1b 2B 128K fixie-ai/ultravox-v0_3-llama-3_2-1b
ultravox-v0_4 8B 8K fixie-ai/ultravox-v0_4
ultravox-v0_4_1-llama-3_1-8b 8B 128K fixie-ai/ultravox-v0_4_1-llama-3_1-8b
ultravox-v0_4_1-mistral-nemo 13B 128K fixie-ai/ultravox-v0_4_1-mistral-nemo
ultravox-v0_5-llama-3_2-1b 2B 128K fixie-ai/ultravox-v0_5-llama-3_2-1b
ultravox-v0_5-llama-3_1-8b 8B 128K fixie-ai/ultravox-v0_5-llama-3_1-8b
ultravox-v0_6-llama-3_1-8b 8B 128K fixie-ai/ultravox-v0_6-llama-3_1-8b
ultravox-v0_6-qwen-3-32b 32B 40K fixie-ai/ultravox-v0_6-qwen-3-32b
Table 16: General information about audio modality baseline models.
Model Parameters Context length Hugging Face Hub link Citation
LLaVA-NeXT-Video-7B-hf 7B 4K llava-hf/LLaVA-NeXT-Video-7B-hf Liu et al. (2024c)
Qwen2-VL-2B-Instruct 2B 32K Qwen/Qwen2-VL-2B-Instruct Wang et al. (2024a)
Qwen2-VL-7B-Instruct 7B 32K Qwen/Qwen2-VL-7B-Instruct
Qwen2-VL-72B-Instruct 72B 32K Qwen/Qwen2-VL-72B-Instruct
Qwen2.5-VL-3B-Instruct 3B 128K Qwen/Qwen2.5-VL-3B-Instruct Bai et al. (2025)
Qwen2.5-VL-7B-Instruct 7B 128K Qwen/Qwen2.5-VL-7B-Instruct
Qwen2.5-VL-72B-Instruct 72B 128K Qwen/Qwen2.5-VL-72B-Instruct
Qwen3-VL-2B-Instruct 2B 262K Qwen/Qwen3-VL-2B-Instruct Yang et al. (2025)
Qwen3-VL-8B-Instruct 9B 262K Qwen/Qwen3-VL-8B-Instruct
Qwen2.5-Omni-3B 6B 32K Qwen/Qwen2.5-Omni-3B Xu et al. (2025)
Qwen2.5-Omni-7B 11B 32K Qwen/Qwen2.5-Omni-7B
MiniCPM-o-2_6 9B 32K openbmb/MiniCPM-o-2_6 Yao et al. (2024)
InternVL3_5-4B 5B 40K OpenGVLab/InternVL3_5-4B Wang et al. (2025b)
InternVL3_5-2B-Instruct 2B 40K OpenGVLab/InternVL3_5-2B-Instruct
InternVL3-9B-Instruct 9B 32K OpenGVLab/InternVL3-9B Zhu et al. (2025)
InternVL3-9B 9B 32K OpenGVLab/InternVL3-9B Zhu et al. (2025)
InternVL3-2B 1B 40K OpenGVLab/InternVL3-2B
Table 17: General information about video modality baseline models.
Model Total LabTabVQA RealVQA ruCLEVR ruCommonVQA ruHHH-Image* ruMathVQA
Human Baseline 0.80 0.91 0.63 0.96 0.84 0.89 0.95
Qwen3-Omni-30B-A3B-Inst 0.55 0.56 / 0.62 0.31 / 0.67 0.50 / 0.63 0.56 / 0.86 0.45 / 0.57 0.01 / 0.30
GPT 4 0.48 0.29 / 0.29 0.23 / 0.70 0.31 / 0.53 0.27 / 0.85 0.39 / 0.49 0.01 / 0.17
Qwen2.5-VL-72B-Inst 0.41 0.53 / 0.60 0.21 / 0.65 0.40 / 0.58 0.43 / 0.80 0.35 / 0.44 0.02 / 0.23
Qwen2-VL-72B-Inst 0.33 0.12 / 0.14 0.10 / 0.49 0.29 / 0.55 0.14 / 0.60 0.14 / 0.19 0.01 / 0.06
Qwen2.5-VL-32B-Inst 0.31 0.47 / 0.57 0.16 / 0.50 0.33 / 0.58 0.30 / 0.76 0.26 / 0.40 0.00 / 0.12
Qwen2.5-VL-7B-Inst 0.26 0.19 / 0.27 0.11 / 0.32 0.23 / 0.34 0.33 / 0.56 0.20 / 0.29 0.02 / 0.10
llava-next-110b-hf 0.24 0.06 / 0.10 0.09 / 0.29 0.11 / 0.20 0.31 / 0.58 0.12 / 0.18 0.01 / 0.01
Phi-3.5-vision-inst 0.23 0.12 / 0.18 0.03 / 0.10 0.08 / 0.16 0.21 / 0.40 0.19 / 0.26 0.01 / 0.02
llava-next-72b-hf 0.23 0.09 / 0.11 0.07 / 0.29 0.09 / 0.19 0.21 / 0.49 0.11 / 0.15 0.00 / 0.00
Qwen2.5-Omni-7B 0.23 0.11 / 0.18 0.08 / 0.35 0.13 / 0.32 0.23 / 0.53 0.14 / 0.23 0.03 / 0.08
Qwen2-VL-7B-Inst 0.20 0.01 / 0.13 0.05 / 0.36 0.10 / 0.34 0.19 / 0.50 0.02 / 0.16 0.01 / 0.02
SmolVLM-Inst 0.19 0.01 / 0.16 0.03 / 0.23 0.11 / 0.34 0.14 / 0.52 0.14 / 0.22 0.02 / 0.04
Qwen3-VL-8B-Inst 0.19 0.00 / 0.07 0.11 / 0.45 0.14 / 0.33 0.23 / 0.59 0.05 / 0.15 0.04 / 0.10
Phi-4-multimodal-inst 0.18 0.26 / 0.35 0.11 / 0.30 0.09 / 0.20 0.07 / 0.36 0.04 / 0.08 0.04 / 0.07
MiniCPM-o-2_6 0.18 0.04 / 0.06 0.08 / 0.31 0.15 / 0.33 0.26 / 0.50 0.06 / 0.09 0.00 / 0.00
Qwen2.5-Omni-3B 0.18 0.04 / 0.14 0.07 / 0.25 0.16 / 0.29 0.22 / 0.48 0.04 / 0.21 0.01 / 0.04
InternVL3-9B 0.17 0.01 / 0.04 0.04 / 0.27 0.19 / 0.37 0.09 / 0.45 0.00 / 0.07 0.01 / 0.02
Qwen2-VL-2B-Inst 0.17 0.00 / 0.09 0.06 / 0.28 0.11 / 0.32 0.12 / 0.44 0.03 / 0.09 0.04 / 0.05
gemma-3-27b-it 0.15 0.00 / 0.01 0.05 / 0.38 0.04 / 0.17 0.07 / 0.44 0.00 / 0.02 0.01 / 0.04
granite-vision-3.3-2b 0.14 0.00 / 0.08 0.01 / 0.10 0.00 / 0.08 0.11 / 0.20 0.08 / 0.20 0.02 / 0.02
Qwen2.5-VL-3B-Inst 0.14 0.02 / 0.11 0.03 / 0.13 0.13 / 0.26 0.15 / 0.32 0.03 / 0.18 0.00 / 0.04
Qwen3-VL-2B-Inst 0.12 0.00 / 0.04 0.06 / 0.35 0.11 / 0.27 0.14 / 0.47 0.05 / 0.08 0.01 / 0.04
llava-1.5-13b-hf 0.12 0.01 / 0.16 0.01 / 0.10 0.01 / 0.08 0.10 / 0.33 0.02 / 0.26 0.01 / 0.01
Table 18: Image modality evaluation results (6 tasks out of 11). All tasks metrics are Exact Match / Judge Score. For ruHHH-Image dataset the metrics are Group Exact Match / Group Judge Score. For Human Baseline aggregated results are provided (average of EM and JudgeScore).
Model Total ruNaturalScienceVQA SchoolScienceVQA UniScienceVQA WEIRD ruTiE-Image
Human Baseline 0.80 0.99 0.82 0.13 0.85 0.77
Qwen3-Omni-30B-A3B-Inst 0.55 0.77 / 0.85 0.64 / 0.70 0.11 / 0.34 0.70 / 0.78 0.63 / 0.64
GPT 4 0.48 0.64 / 0.69 0.59 / 0.66 0.10 / 0.40 0.69 / 0.79 0.70 / 0.73
Qwen2.5-VL-72B-Inst 0.41 0.01 / 0.05 0.24 / 0.32 0.11 / 0.25 0.65 / 0.76 0.63 / 0.69
Qwen2-VL-72B-Inst 0.33 0.25 / 0.40 0.54 / 0.68 0.18 / 0.33 0.31 / 0.48 0.65 / 0.69
Qwen2.5-VL-32B-Inst 0.31 0.00 / 0.03 0.00 / 0.04 0.01 / 0.09 0.47 / 0.77 0.43 / 0.62
Qwen2.5-VL-7B-Inst 0.26 0.13 / 0.17 0.16 / 0.32 0.06 / 0.13 0.36 / 0.62 0.24 / 0.47
llava-next-110b-hf 0.24 0.25 / 0.34 0.25 / 0.45 0.06 / 0.13 0.24 / 0.33 0.54 / 0.55
Phi-3.5-vision-inst 0.23 0.32 / 0.53 0.22 / 0.33 0.07 / 0.11 0.41 / 0.60 0.33 / 0.35
llava-next-72b-hf 0.23 0.19 / 0.27 0.40 / 0.47 0.07 / 0.13 0.27 / 0.33 0.52 / 0.55
Qwen2.5-Omni-7B 0.23 0.09 / 0.23 0.12 / 0.24 0.06 / 0.12 0.30 / 0.60 0.28 / 0.48
Qwen2-VL-7B-Inst 0.20 0.04 / 0.38 0.03 / 0.52 0.09 / 0.19 0.08 / 0.38 0.11 / 0.58
SmolVLM-Inst 0.19 0.19 / 0.36 0.11 / 0.27 0.06 / 0.10 0.36 / 0.44 0.14 / 0.24
Qwen3-VL-8B-Inst 0.19 0.03 / 0.08 0.14 / 0.20 0.10 / 0.22 0.08 / 0.17 0.39 / 0.41
Phi-4-multimodal-inst 0.18 0.05 / 0.12 0.30 / 0.40 0.08 / 0.15 0.02 / 0.07 0.44 / 0.46
MiniCPM-o-2_6 0.18 0.07 / 0.19 0.20 / 0.35 0.01 / 0.10 0.20 / 0.24 0.34 / 0.41
Qwen2.5-Omni-3B 0.18 0.07 / 0.23 0.09 / 0.24 0.06 / 0.12 0.14 / 0.60 0.13 / 0.36
InternVL3-9B 0.17 0.17 / 0.25 0.37 / 0.44 0.06 / 0.16 0.01 / 0.16 0.21 / 0.40
Qwen2-VL-2B-Inst 0.17 0.04 / 0.32 0.04 / 0.38 0.08 / 0.14 0.09 / 0.31 0.16 / 0.44
gemma-3-27b-it 0.15 0.05 / 0.12 0.49 / 0.55 0.04 / 0.28 0.00 / 0.12 0.21 / 0.25
granite-vision-3.3-2b 0.14 0.27 / 0.38 0.08 / 0.22 0.04 / 0.07 0.20 / 0.52 0.20 / 0.26
Qwen2.5-VL-3B-Inst 0.14 0.02 / 0.26 0.04 / 0.22 0.05 / 0.11 0.09 / 0.50 0.08 / 0.35
Qwen3-VL-2B-Inst 0.12 0.01 / 0.04 0.14 / 0.19 0.05 / 0.11 0.03 / 0.10 0.21 / 0.25
llava-1.5-13b-hf 0.12 0.01 / 0.37 0.01 / 0.25 0.00 / 0.09 0.01 / 0.49 0.03 / 0.22
Table 19: Image modality evaluation results (5 tasks out of 11). All tasks metrics are Exact Match / Judge Score. For UniScienceVQA the Human Baseline is crowd, not an expert. For Human Baseline aggregated results are provided (average of EM and JudgeScore).
Model Total AQUARIA ruEnvAQA ruTiE-Audio ruSLUn
Human Baseline 0.895 0.98 0.95 0.75 0.91
Qwen3-Omni-30B-A3B-Instruct 0.56 0.69 / 0.77 0.70 / 0.78 0.43 / 0.44 0.39 / 0.28
Qwen2.5-Omni-7B 0.47 0.55 / 0.67 0.55 / 0.70 0.36 / 0.41 0.37 / 0.18
Qwen2.5-Omni-3B 0.38 0.41 / 0.58 0.36 / 0.64 0.30 / 0.36 0.35 / 0.04
MiniCPM-o-2_6 0.37 0.40 / 0.55 0.47 / 0.64 0.25 / 0.32 0.31 / 0.01
ultravox-v0_6-qwen-3-32b 0.32 0.39 / 0.42 0.37 / 0.40 0.47 / 0.51 0.00 / 0.00
ultravox-v0_5-llama-3_1-8b 0.31 0.30 / 0.37 0.34 / 0.41 0.29 / 0.31 0.31 / 0.15
ultravox-v0_4_1-llama-3_1-8b 0.31 0.29 / 0.36 0.35 / 0.42 0.31 / 0.32 0.28 / 0.11
ultravox-v0_4 0.30 0.30 / 0.38 0.34 / 0.43 0.29 / 0.31 0.30 / 0.09
ultravox-v0_6-llama-3_1-8b 0.30 0.29 / 0.36 0.26 / 0.40 0.30 / 0.32 0.31 / 0.16
ultravox-v0_3 0.28 0.29 / 0.35 0.35 / 0.43 0.27 / 0.29 0.22 / 0.04
ultravox-v0_4_1-mistral-nemo 0.26 0.18 / 0.34 0.27 / 0.42 0.18 / 0.34 0.26 / 0.14
audio-flamingo-3-hf 0.26 0.21 / 0.42 0.41 / 0.59 0.15 / 0.28 0.00 / 0.00
Qwen2-Audio-7B-Instruct 0.22 0.18 / 0.43 0.15 / 0.52 0.17 / 0.28 0.05 / 0.00
ultravox-v0_2 0.14 0.01 / 0.25 0.00 / 0.38 0.01 / 0.25 0.23 / 0.00
ultravox-v0_5-llama-3_2-1b 0.12 0.04 / 0.25 0.08 / 0.29 0.05 / 0.23 0.00 / 0.00
Qwen-Audio-Chat 0.12 0.01 / 0.28 0.01 / 0.40 0.01 / 0.24 0.00 / 0.00
ultravox-v0_3-llama-3_2-1b 0.12 0.06 / 0.24 0.08 / 0.29 0.05 / 0.20 0.00 / 0.00
SeaLLMs-Audio-7B 0.10 0.08 / 0.14 0.02 / 0.20 0.14 / 0.23 0.00 / 0.01
Table 20: Audio modality evaluation results. All tasks metrics are Exact Match / Judge Score. For ruSLUn dataset the metrics are Intent Exact Match / Slot F1 Score. For Human Baseline aggregated results are provided (average of EM and JudgeScore).
Model Total CommonVideoQA RealVideoQA ruHHH-Video
Human Baseline 0.92 0.96 0.96 0.84
Qwen2.5-VL-72B-Instruct 0.63 0.57 / 0.64 0.65 / 0.73 0.53 / 0.63
Qwen3-VL-8B-Instruct 0.58 0.53 / 0.60 0.61 / 0.69 0.46 / 0.56
Qwen2-VL-72B-Instruct 0.56 0.49 / 0.62 0.57 / 0.70 0.34 / 0.62
Qwen2.5-VL-7B-Instruct 0.52 0.49 / 0.56 0.58 / 0.67 0.41 / 0.44
Qwen2.5-Omni-7B 0.44 0.42 / 0.55 0.44 / 0.61 0.24 / 0.39
Qwen2.5-VL-3B-Instruct 0.43 0.39 / 0.49 0.47 / 0.60 0.25 / 0.37
Qwen3-VL-2B-Instruct 0.42 0.41 / 0.48 0.51 / 0.59 0.20 / 0.31
Qwen3-Omni-30B-A3B-Instruct 0.41 0.00 / 0.00 0.64 / 0.72 0.48 / 0.63
MiniCPM-o-2_6 0.37 0.33 / 0.49 0.41 / 0.57 0.13 / 0.30
Qwen2.5-Omni-3B 0.34 0.30 / 0.46 0.33 / 0.54 0.12 / 0.27
InternVL3-9B-Instruct 0.32 0.27 / 0.30 0.25 / 0.32 0.33 / 0.44
InternVL3-9B 0.32 0.28 / 0.31 0.26 / 0.31 0.32 / 0.42
Qwen2-VL-7B-Instruct 0.30 0.09 / 0.50 0.12 / 0.61 0.04 / 0.45
InternVL3_5-4B 0.29 0.27 / 0.31 0.28 / 0.32 0.24 / 0.32
LLaVA-NeXT-Video-7B-hf 0.13 0.09 / 0.21 0.09 / 0.24 0.03 / 0.09
Table 21: Video modality evaluation results. All tasks metrics are Exact Match / Judge Score. For ruHHH-Video dataset the metrics are Group Exact Match / Group Judge Score. For Human Baseline aggregated results are provided (average of EM and JudgeScore).
The exact results of the finished submissions of the models from Table 15 are presented in Tables 18, 19. The results of the models from Table 16 are presented in Table 20. The evaluation results of the models from Table 17 are stated in Table 21.
Evaluated on vision (image) modality models are strongest on natural-image semantics — object/scene understanding, object functions, and everyday knowledge (RealVQA, ruCommonVQA, WEIRD, ruTiE-Image) and show decent spatial relations and multi-object reasoning (RealVQA, ruCLEVR). Performance drops on diagrammatic/scientific QA with OCR (UniScienceVQA, ruNaturalScienceVQA) and tables (LabTabVQA), exposing gaps in text extraction, scheme recognition, and cell-level grounding. Math (ruMathVQA) is brittle: EM lags JS, indicating partially correct solutions that miss the required final answer. Harder compositional and counterfactual reasoning (RealVQA, ruCommonVQA) still separates model tiers, especially smaller checkpoints. Overall, weaknesses concentrate in OCR/diagram/table parsing and structured compositional reasoning, while strengths lie in natural-image commonsense and basic spatial reasoning.
Moving to audio modality evaluations, the relative strengths cluster around acoustic scene understanding and temporal/comparative reasoning over environmental audio (ruEnvAQA), with partial competence in social/interaction cues and topic/scene grounding (AQUARIA, ruTiE-Audio) captured by higher JS. Weaknesses are pronounced in speech recognition (RuSLUn, ruTiE-Audio— EM), speaker/turn handling (diarization), and format-faithful final answer extraction. Improving ASR robustness, speaker attribution, and constrained decoding or answer templates should convert many high-JS outputs into EM gains.
As for the video modality evaluations, the current relative strengths lie in scene/object recognition and short event recognition. Clear weaknesses persist in temporal perception/localization, action-sequence reasoning, mutual object localization, counting, and cause-and-effect—skills central to CommonVideoQA/RealVideoQA. Low ruHHH-Video scores further reveal brittleness on ethics-aware interpretation and social context. Practically, improving temporal grounding (longer context windows or frame selection), causal reasoning, and answer-format control (constrained decoding/final-answer extraction) should convert many near-misses (high JS, low EM) into measurable EM gains.
F.2 Human Baseline Details
Human baseline values were obtained by evaluating the aggregate responses of annotators on control tasks. Prior to metric calculation, we conducted an additional identification of annotators who performed labeling tasks with low quality. To identify such annotators during the labeling process, we employed control tasks with automated correctness checking; incorrect answers reduced the annotator’s skill score. During post-processing, we filtered out annotators who consistently made errors and whose responses did not align with the majority vote across all their submissions.
To evaluate the average human level we collected crowd-source [^25] Annotation provided by ABC Elementary platform https://app.elementary.center.
For expert tasks, in addition to the basic human baseline, we also performed expert annotation. We aggregate experts’ answers with an overlap of 3 for expert human baseline scores  [^26] The observed quantity is a consequence of the limited pool of domain experts in niche specializations, who also had no prior involvement in the annotation of the datasets.. The experts annotating answers for the human baseline had not been involved in the original dataset creation.
Dataset Overlap Num samples Total, \$ Per item, \$ Per hour, \$ IAA
AQUARIA 5 786 324.42 0.41 7.38 93.87%
CommonVideoQA 5 1200 2364.12 1.97 8.53 92.41%
LabTabVQA 5 339 518.99 0.31 11.04 87.91%
RealVQA 5 1 010 405.82 0.40 8.54 69.65%
RealVideoQA 5 671 785.54 1.17 8.53 92.19%
ruCLEVR 5 2 063 1 440.15 0.70 7.40 93.36%
ruCommonVQA 5 2 922 10 401.71 3.56 8.54 79.24%
ruEnvAQA 5 644 239.58 0.37 7.38 89.46%
ruHHH-Image 5 610 276.94 0.45 8.54 90.25%
ruHHH-Video 5 911 638.40 0.70 708 91.28%
ruMathVQA (crowd) 5 2 975 214.71 0.07 1.10 85.01%
ruMathVQA (expert) 5 2 975 1 363.58 2.29 1.35 83.09%
ruNaturalScienceVQA (crowd) 5 403 119.55 0.30 6.09 90.37%
ruNaturalScienceVQA (expert) 3 403 123.08 0.31 9.65 96.69%
RuSLUn 5 741 133.00 0.03 2.88 81.05%
ruTiE-Audio 3 500 65.16 0.13 3.62 81.00%
ruTiE-Image 3 500 65.16 0.13 3.62 85.33%
SchoolScienceVQA (crowd) 5 1 750 1 293.83 0.74 8.54 67.56%
SchoolScienceVQA (expert) 3 1 750 1 270.70 0.73 8.94 81.23%
UniScienceVQA 5 1 150 102.4 0.09 2.41 45.19%
WEIRD 5 889 109.1 0.10 9.4 90.84%
Table 22: Payrates and total expenses for human baseline annotation.
The ABC Elementary annotation platform ensures the necessary data anonymity during processing. The hourly compensation offered is above the minimum wage per hour in Russia (see Table 22). Annotators are made aware of potentially sensitive topics within the data, including politics, societal minorities, and religion. The data collection process undergoes a mandatory quality evaluation, featuring an automated annotation quality check through honeypot tasks.